### v1

# A Versatility Measure for Parametric Risk Models

Abstract
Parametric statistical methods play a central role in analyzing risk through its underlying frequency and severity components. Given the wide availability of numerical algorithms and high-speed computers, researchers and practitioners often model these separate (although possibly statistically dependent) random variables by fitting a large number of parametric probability distributions to historical data and then comparing goodness-of-fit statistics. However, this approach is highly susceptible to problems of overfitting because it gives insufficient weight to fundamental considerations of functional simplicity and adaptability. To address this shortcoming, we propose a formal mathematical measure for assessing the versatility of frequency and severity distributions prior to their application. We then illustrate this approach by computing and comparing values of the versatility measure for a variety of probability distributions commonly used in risk analysis.
Keywords: Parametric probability distributions; risk models; complexity; versatility.
1 Introduction
Parametric statistical methods play a central role in analyzing risk through its underlying frequency and severity components. [^1] See, e.g., Klugman, Panjer, and Willmot (2019). Frequencies, which represent counts of the number of damage-causing events to occur within a specified time period, generally are modeled by nonnegative discrete random variables, $X\in\mathbb{Z}_{0}^{+}$, whereas severities, which measure the individual damage amounts associated with these events, typically are modeled by nonnegative continuous random variables, $Y\in\mathbb{R}_{0}^{+}$ (which may be statistically dependent on $X$). [^2] The terms “frequency” and “severity” are routinely used in the fields of operational risk management and actuarial finance, but various alternatives (such as “rate”, “probability”, or “likelihood” for frequency and “intensity”, “impact”, or “consequence” for severity) often appear in other scientific disciplines (epidemiology, environmental science, etc.). Given the broad availability of efficient numerical algorithms and high-speed computers, researchers and practitioners often select a parametric frequency or severity distribution for estimation purposes by fitting a large number of potential distributions to relevant historical data and then comparing the associated goodness-of-fit statistics. Unfortunately, this approach can lead to model overfitting.
In the context of modeling frequency and severity data, the most obvious type of overfitting involves selecting a distribution function with an excessively large number of parameters, $k$, compared to the number of observations, $n$. Such “parametric overfitting” is analogous to constructing models with too many explanatory variables for a given sample size in a conventional linear-regression setting, and can be addressed in the same way; that is, by employing goodness-of-fit statistics and/or tests that impose explicit penalties for each parameter estimated, essentially balancing the costs and benefits of increasing $k$ for a given $n$.
Another, less conspicuous type of overfitting that can occur when modeling frequencies and severities is to select distributions whose functional forms are characterized by idiosyncrasy (i.e., excessive complexity and/or contrivance of structure). Complexity is analogous to including linear-regression covariates that are redundant (i.e., linearly dependent on some subset of other explanatory variables), whereas contrivance is analogous to including covariates that possess no plausible relationship with the dependent variable. In linear regression, the most straightforward means of avoiding redundant explanatory variables is to exclude them through multicollinearity analysis. For irrelevant explanatory variables, the solution is even simpler: they may be discarded because of the absence of any connection to the dependent variable. [^3] This is justified by the scientific method, to be discussed in Section 2, which generally does not support the inclusion of factors based on empirical statistical relationships without some theoretical connection. Unfortunately, there currently is no clear defense against idiosyncratic distributions in risk analysis because no distribution functions defined on $x\in\mathbb{Z}_{0}^{+}$ and $y\in\mathbb{R}_{0}^{+}$are manifestly inappropriate, a priori, for frequencies and severities, respectively.
In the present investigation, we propose a novel approach to mitigating this type of “idiosyncratic overfitting” based on the idea that, for a fixed number of parameters, probability distributions used to model frequencies and severities must demonstrate a reasonable degree of functional versatility (construed as the opposite of idiosyncrasy). We begin, in Section 2, by developing the concept of versatility based on underlying components of functional simplicity and adaptability, which in turn are motivated by the scientific method. Next, mathematical techniques for measuring simplicity and adaptability are explored in Sections 3 and 4, leading to a tentative measure of functional versatility. This versatility measure is refined and formalized in Section 5, then illustrated through application to a number of common frequency and severity models in Section 6. Finally, in Section 7, we offer some closing observations and conclusions.
2 Model Versatility
The list below provides preliminary definitions of various terms employed in the Introduction. Although somewhat terse, these definitions will be refined (both verbally and mathematically) in the present and subsequent sections.
  • Complexity - The degree of intricacy or disorder in the form of a mathematical function.
  • Contrivance - The degree of artificiality or rigidity in the form of a mathematical function.
  • Idiosyncrasy - The total amount of complexity and/or contrivance associated with the form of a mathematical function.
  • Simplicity - The degree of plainness or coherence in the form of a mathematical function (taken as the opposite of complexity).
  • Adaptability - The degree of naturalness or resiliency in the form of a mathematical function (taken as the opposite of contrivance).
  • Versatility - The overall level of simplicity and/or adaptability associated with the form of a mathematical function (taken as the opposite of idiosyncrasy).
Clearly, the first three terms describe potential “negative” characteristics of parametric models, undesirable because they indicate the presence of unnecessary complications or restrictions in the modeling process. The last three terms, which represent the “positive” opposites of the first three, thus serve as appropriate objectives of model selection. Although it might seem attractive to shorten the list by referring to the positive concepts in terms of their negative counterparts (e.g., replacing simplicity by “inverse-complexity”) or vice versa, this would create expositional difficulties. In particular, it is natural to use idiosyncrasy to describe the model-selection problems created by its components (complexity and contrivance), but to connect these issues to the scientific method through the corresponding components of versatility (simplicity and adaptability), as shown in the following subsection.
2.1 The Scientific Method
In both the natural and social sciences, the scientific method constitutes the generally accepted framework for guiding and marking the progress of human knowledge. Despite lacking a formal definition and possessing minor variations across research fields, this method typically includes the following steps: [^4] See, e.g., Powers (2012).
(I) The statement of a null hypothesis ($H_{0}$) describing a scientific relationship among observable quantities. This hypothesis, based on derivations from theory [^5] As noted in the Introduction, the scientific method generally requires a theoretical basis for hypotheses. One motivation for this is to avoid giving formal “scientific” recognition to statistically significant results in experimental studies of paranormal phenomena (whose causal explanations lie outside currently accepted science). However, certain mainstream fields (e.g., medical research) routinely overlook this requirement for pragmatic reasons, and thus satisfy a somewhat lower standard of knowledge. (For example, experimental studies may demonstrate the efficacy of a particular drug in treating a certain disease without a clear understanding of the physiological mechanism involved.) (usually guided by inferences from previous observations), summarizes the scientist’s best understanding of the indicated relationship in the simplest way possible.
(II) The identification of an observable phenomenon that is predicted by the underlying hypothesis and not otherwise explicable.
(III) The design of an experiment comprising a reasonably large number of individual observations, each of which reflects the presence/absence of the identified phenomenon. This includes such critical issues as identifying the statistical properties of the observations, how many observations will be made, and the level of significance (i.e., probability of Type 1 error, $\pi_{1}$), for rejecting the null hypothesis.
(IV) The execution and subsequent replication of the experiment described in (III) by numerous independent researchers. This both reduces the possibility of systematic errors associated with individual researchers and enhances the statistical significance of results by adding more observations.
Clearly, the above outline draws heavily from the paradigm of statistical hypothesis testing. In fact, steps (I) through (III) essentially form an application of hypothesis testing to a specific scientific problem, accomplished by writing the scientific relationship under investigation in the formal mathematical terms of a null hypothesis. For the present study, we will view the frequency- and severity-modeling problem as an application of the scientific method in which the null hypothesis takes the form
$H_{0}:\>W\sim F_{0}\equiv F_{W\mid\boldsymbol{a}}^{\left(0\right)}\left(w\right),$ (1)
where $F_{0}\equiv F_{W\mid\boldsymbol{\alpha}}^{\left(0\right)}\left(w\right)$ represents the cumulative distribution function of either $W=X\in\mathbb{\mathbb{Z}}_{0}^{+}$ or $W=Y\in\mathbb{R}_{0}^{+}$, with unknown $k$-dimensional parameter vector $\boldsymbol{a}=\left[a_{1},\ldots,a_{k}\right]$. The following subsections address characteristics of $F_{0}$ that are particularly desirable in this setting.
2.2 Parsimony and Functional Simplicity
As can be seen from step (I) of the outline, the scientific method incorporates a bias in favor of forming the simplest feasible null hypothesis consistent with theoretical derivations and/or empirical observations. This approach reflects the principle of parsimony often associated with Occam’s Razor: “Entities are not to be multiplied without necessity.” [^6] Although attributed to William of Ockham (1287-1347), the first published statement of this dictum (in Latin) was provided by the Irish Franciscan friar and philosopher, John Punch, in 1639. (which is sometimes rendered in modern English as “The simplest explanation is usually the best.”). However, this principle was known in the West at least as early as Aristotle (384-322 BC), who wrote (in his Posterior Analytics): “We may assume the superiority, other things being equal, of the demonstration which derives from fewer postulates or hypotheses.”
The principle of parsimony is embedded in many statistical model-selection procedures through the use of goodness-of-fit statistics (such as the Akaike and Bayesian information criteria) and/or hypothesis-testing procedures (such as minimum Chi-squared) that include explicit penalties for employing additional parameters. As observed in the Introduction, such methods can address the problem of overfitting through over-parameterized probability distributions, but not that of overfitting through idiosyncratic (i.e., excessively complex and/or contrived) distribution functions.
To clarify these relationships, let the real-valued quantities $\mathcal{I}\left(F_{0}\right)\geq 0$, $\mathcal{C}\left(F_{0}\right)\geq 0,$ and $\mathcal{T}\left(F_{0}\right)\geq 0$ denote hypothetical heuristic measures of the intrinsic idiosyncrasy, complexity, and contrivance of the distribution function $F_{0}$, where each quantity is assumed to be independent of the number of parameters, $k$. In addition, set $\mathcal{V}\left(F_{0}\right)=1/\mathcal{I}\left(F_{0}\right),$ $\mathcal{S}\left(F_{0}\right)=1/\mathcal{C}\left(F_{0}\right),$ and $\mathcal{A}\left(F_{0}\right)=1/\mathcal{T}\left(F_{0}\right)$ as the corresponding inverse (i.e., opposite) quantities: versatility, simplicity, and adaptability, respectively. Then, from the discussion of the preceding paragraph, it follows (heuristically) that
$\mathcal{I}\left(F_{0}\right)=\mathcal{I}_{1}\left(\mathcal{C}^{\left(+\right)},\mathcal{T}^{\left(+\right)}\right)$
$=\mathcal{I}_{2}\left(\mathcal{S}^{\left(-\right)},\mathcal{A}^{\left(-\right)}\right),$ (2)
where the superscripts indicate the direction of the argument’s relationship with $\mathcal{I}$ (i.e., $\left(+\right)=\textrm{positive}$ or $\left(-\right)=\textrm{negative}$).
Now let $\mathcal{P}\left(F_{0}\right)\geq 0$ denote a hypothetical real-valued measure of the parsimony of $F_{0}$, taken to be negatively related to both its number of parameters and functional complexity (because both quantities oppose the objective of simplicity), but independent of functional contrivance (because unnaturally manipulated functions can possess simple or complex structures). We then can write (heuristically)
$\mathcal{P}\left(F_{0}\right)=\mathcal{P}_{1}\left(k^{\left(-\right)},\mathcal{C}^{\left(-\right)}\right)$
$=\mathcal{P}_{2}\left(k^{\left(-\right)},\mathcal{S}^{\left(+\right)}\right).$ (3)
Functional simplicity, $\mathcal{S}\left(F_{0}\right)$, which both increases parsimony and reduces idiosyncrasy, will be investigated further in Section 3.
2.3 Robustness and Functional Adaptability
In addition to parsimony, the scientific method also entails a bias in favor of the status quo, operationalized through the choice of $\pi_{1}$ in step (III) of the above outline. By fixing the probability of Type 1 error, $\pi_{1}=\Pr\left\{\textrm{Reject }H_{0}\mid H_{0}\textrm{ is true}\right\}$ at a relatively low level, while not imposing any maximum value on the probability of Type 2 error, $\pi_{2}=\Pr\left\{\textrm{Retain }H_{0}\mid H_{0}\textrm{ is false}\right\}$, the paradigm embodies what is sometimes called the Principle of Laplace: “The weight of the evidence should be proportioned to the strangeness of the facts.” [^7] The original statement by Pierre-Simon Laplace (1749-1827) – “[W]e ought to examine [inexplicable phenomena] with an attention all the more scrupulous as it appears more difficult to admit them.” (in French) – appeared in his “Philosophical Essay on Probabilities” (1814). The “Principle” is a paraphrase of this admonition (in French) first published by Swiss professor of psychology, Théodore Flournoy, in 1899. That is, if we currently believe the null hypothesis to be true, then we should not abandon this belief unless confronted by compelling evidence; and the greater the deviation in belief from the null hypothesis, the more substantial the evidence necessary to justify it. Applying this concept to the frequency- and severity-modeling problem, one can see that it suggests the probability distribution under consideration manifest robustness in the sense of serving as a reasonable model for a relatively large collection of possible data sets (and thus being more resistant to rejection under hypothesis testing).
Letting $\mathcal{R}\left(F_{0}\right)\geq 0$ denote a hypothetical real-valued measure of the robustness of $F_{0}$, we will take this quantity to be positively related to its number of parameters (more of which permit greater accuracy of fit), negatively related to its degree of functional contrivance (because unnaturally manipulated functions are less adaptable to a wide range of empirical observations), and independent of functional complexity (because unnaturally complicated functions can enjoy various degrees of adaptability). We thus write (again heuristically)
$\mathcal{R}\left(F_{0}\right)=\mathcal{R}_{1}\left(k^{\left(+\right)},\mathcal{T}^{\left(-\right)}\right)$
$=\mathcal{R}_{2}\left(k^{\left(+\right)},\mathcal{A}^{\left(+\right)}\right).$ (4)
Functional adaptability, $\mathcal{A}\left(F_{0}\right)$, which both increases robustness and reduces idiosyncrasy, will be studied in Section 4.
2.4 Functional Versatility
Since both parsimony and robustness are negatively related to idiosyncrasy, it might seem that seeking distribution functions, $F_{0}$, with higher levels of these characteristics is an effective means of avoiding functions that are idiosyncratic. However, as indicated by equations (3) and (4), these quantities may depend explicitly on $k$, in which case they can be increased simply by adjusting the number of parameters, rather than by increasing functional simplicity and adaptability (the quantities directly relevant to idiosyncrasy).
With this in mind, we treat functional versatility as the inverse of idiosyncrasy:
$\mathcal{V}\left(F_{0}\right)=\mathcal{V}\left(\mathcal{S}^{\left(+\right)},\mathcal{A}^{\left(+\right)}\right),$ (5)
and propose a formal expression for this quantity in Section 5.
The following table provides a summary of the various specialized concepts employed in the present subsection. From this table, it is clear that the distinctions between parsimony and simplicity, on the one hand, and robustness and adaptability, on the other, are fairly minor. However, it is important to differentiate between the concepts in each pair to recognize the significant role played by the number of distribution parameters.
Table 1. Concepts Associated with Avoiding Model Overfitting
Concept Heuristic Measure Inverse
Simplicity $\mathcal{S}\left(F_{0}\right)=1/\mathcal{C}\left(F_{0}\right)$ Complexity
Parsimony $\mathcal{P}\left(F_{0}\right)=\mathcal{P}_{2}\left(\mathcal{S}^{\left(+\right)},k^{\left(-\right)}\right)$
$=\mathcal{P}_{1}\left(\mathcal{C}^{\left(-\right)},k^{\left(-\right)}\right)$
Adaptability $\mathcal{A}\left(F_{0}\right)=1/\mathcal{T}\left(F_{0}\right)$ Contrivance
Robustness $\mathcal{R}\left(F_{0}\right)=\mathcal{R}_{2}\left(\mathcal{A}^{\left(+\right)},k^{\left(+\right)}\right)$
$=\mathcal{R}_{1}\left(\mathcal{T}^{\left(-\right)},k^{\left(+\right)}\right)$
Versatility $\mathcal{V}\left(F_{0}\right)=\mathcal{V}\left(\mathcal{S}^{\left(+\right)},\mathcal{A}^{\left(+\right)}\right)$ Idiosyncrasy
$=1/\mathcal{I}\left(F_{0}\right)$
3 Measuring Simplicity
3.1 Two Types of Complexity
Recognizing that functional simplicity is the opposite of functional complexity, it seems reasonable to measure the former by selecting and inverting some measure of the latter, many of which have been proposed in the research literatures of various fields: statistical thermodynamics, information theory, algorithmic analysis, etc. Although similarities exist among certain of these measures (especially those based on some form of “entropy”), differences in terminology and perspective necessitate a precise identification of the specific type of complexity we wish to model. In particular, we must determine which of two distinct forms of complexity is more appropriate for assessing probability mass functions (PMFs) and probability density functions (PDFs) in the present context: Type A, or notational complexity, based on the number and arrangement of mathematical symbols and operations required to express the function; and Type B, or morphological complexity, based on the function’s geometric/topological properties as a figure plotted in a two-dimensional coordinate system.
With regard to Type A, it is important to note that the complexity of a PMF/PDF’s notational form should not be conflated with notions of artificial (and possibly tendentious) design, even if we know or suspect a particular PMF/PDF to have been constructed by a clever probabilist with specific objectives in mind. Certainly, functions can be devised to achieve high levels of Kolmogorov (algorithmic) complexity (also called Kolmogorov entropy), meaning that they cannot be compressed into short string lengths in a mathematical coding system. [^8] See Li and Vitányi (2019) for a recent treatment of Kolmogorov complexity/entropy. However, such constructions necessitate intentionally injecting elements of disorder or disorganization, since functions formed according to an ordered plan or pattern (even if somewhat complicated) possess less Kolmogorov complexity because they are more compressible.
In the case of Type B, we encounter an apparent paradox: PMFs/PDFs with greater Shannon (information-theoretic) entropy (also called Shannon information) tend to possess simpler, flatter shapes. [^9] See Gray (2011) for a comprehensive overview of Shannon entropy. For example, the maximum-entropy discrete distribution on $x\in\mathbb{\mathbb{Z}}_{0}^{+}$, subject to a fixed mean, is given by the $\textrm{Geometric}\left(\varrho\right)$ PMF ($f_{X\mid\varrho}\left(x\right)=\varrho\left(1-\varrho\right)^{x}$, for $\varrho\in\left(0,1\right)$), and the maximum-entropy continuous distribution on $y\in\mathbb{R}_{0}^{+}$, again subject to a fixed mean, is given by the $\textrm{Exponential}\left(\lambda\right)$ PDF ($f_{Y\mid\lambda}\left(y\right)=\lambda e^{-\lambda y}$, for $\lambda\in\mathbb{R}^{+}$). This is attributable to the fact that, for a given sample space and specified constraints (e.g., a fixed mean or other moment), PMFs/PDFs providing greater Shannon entropy, as measured by the expected “surprise” from a random trial, are those that spread the distribution’s probability more evenly across the sample space. Essentially, we can say (informally, at least) that increasing the disorder or disorganization of random draws is achieved by decreasing the complexity of the distribution function’s shape.
When choosing between the two complexity types, one can see that Type B offers a clear advantage: it is possible to select a simpler PMF/PDF just by favoring greater values of the distribution function’s Shannon entropy,
$\textrm{H}_{F_{0}}\left(W\right)=\textrm{E}_{F_{0}}\left[-\ln\left(f_{0}\left(W\right)\right)\right],$ (6)
without having to analyze and evaluate the mathematical symbols and operations involved in expressing $f_{0}$. As shown in Su et al. (2021), numerous methods for assessing the complexity of functional forms have been proposed in the research literature, but none is broadly accepted, and none offers the analytical tractability of (6).
Another, more substantive, advantage of Type B complexity is demonstrated by following example. Consider the PDF
$f_{Y\mid\delta,\ell}\left(y\right)=K\exp\left(-\dfrac{y}{\delta}\left[\ell+1+\sin\left(\varepsilon\eta_{1}y\right)+\sin\left(\varepsilon\eta_{2}y\right)+\cdots+\sin\left(\varepsilon\eta_{\ell}y\right)\right]\right),$
defined on $y\in\mathbb{R}_{0}^{+}$, where:
$\delta\in\mathbb{R}^{+}$ and $\ell\in\left\{1,2,\ldots\right\}$ are parameters;
$\varepsilon\in\mathbb{R}_{0}^{+}$ and the $\eta_{i}\in\left\{0,1\right\}$ are fixed constants; and
[…]
From this formulation, it is easy to see that each of the two conditions, (i) $\varepsilon=0$ and (ii) $\eta_{i}=0$ for all $i$, yields the maximum-entropy Exponential PDF mentioned above.
Setting $\varepsilon=1$ and fixing the values of the parameters $\delta$ and $\ell$, one can manipulate the Type A complexity of $f_{Y\mid\delta,\ell}\left(y\right)$ by selecting the individual $\eta_{i}$ to generate more or less structured (i.e., more or less Kolmogorov compressible) sequences, $\eta_{1},\eta_{2},\ldots,\eta_{\ell}$. [^10] For example, if $m=10$, then the sequence $0,1,0,1,0,1,0,1,0,1$ is more structured (Kolmogorov compressible) than the sequence $0,1,1,1,0,1,0,0,1,0$. As this occurs, Type B complexity will fluctuate independently of Type A complexity, depending on the sequence $\eta_{1},\eta_{2},\ldots,\eta_{\ell}$ through only the sum ${\textstyle\sum_{i=1}^{\ell}\eta_{i}}$, and achieving its lowest levels when this sum is close to 0 (so that $f_{Y\mid\delta,\ell}\left(y\right)$ approaches the Exponential PDF).
Now consider what happens as $\varepsilon\rightarrow 0^{+}$. For any fixed selection of $\eta_{1},\eta_{2},\ldots,\eta_{\ell}$, Type A complexity remains constant as $\varepsilon$ decreases (because the formula for $f_{Y\mid\delta,\ell}\left(y\right)$ remains essentially the same), whereas Type B complexity decreases (because $f_{Y\mid\delta,\ell}\left(y\right)$ approaches the Exponential PDF). Clearly, this characteristic of Type A is undesirable for our purposes because it fails to differentiate between intrinsically simple PMFs/PDFs (such as those close to the Exponential PDF) and substantially more complex functions.
3.2 Exponential Entropy and $\boldsymbol{p^{\textrm{th}}}$ Power Means
Having concluded that Type B complexity is more appropriate for the current context, we now turn to developing a measure of functional simplicity beginning with the concept of Shannon entropy. Although the intuition underlying (6) is fairly straightforward in the case of discrete random variables, this is less true for continuous random variables, where entropy often is motivated by analogy with the discrete case.
For a frequency $X\sim F_{X}\left(x\right)$ defined on $x\in\mathbb{\mathbb{Z}}_{0}^{+}$, the Shannon entropy is given by
$\textrm{H}_{F_{X}}\left(X\right)=\textrm{E}_{F_{X}}\left[-\ln\left(f_{X}\left(X\right)\right)\right]$
$={\displaystyle\sum_{x=0}^{\infty}}\left(-\ln\left(f_{X}\left(x\right)\right)\right)f_{X}\left(x\right),$ (7)
where $-\ln\left(f_{X}\left(x\right)\right)$ denotes the surprise associated with the random outcome $X=x$. Since $-\ln\left(\cdot\right)$ is a strictly decreasing function, this is consistent with the intuition that outcomes with lower probability generate greater surprise. The formal justification for the logarithmic form is that units of surprise arising from statistically independent observations must be additive.
If the sample space of $X$ were bounded – for example, $x\in\left\{0,1,\ldots,B\right\}$ – then it is quite easy to show that the associated entropy,
$\textrm{H}_{F_{X}}\left(X\right)={\displaystyle\sum_{x=0}^{B}}\left(-\ln\left(f_{X}\left(x\right)\right)\right)f_{X}\left(x\right),$
would be maximized when the PMF is Discrete $\textrm{Discrete Uniform}\left(0,B\right)$; that is, $f_{X\mid B}\left(x\right)=\tfrac{1}{B+1}$ for all $x$ and $\textrm{H}_{F_{X\mid B}}\left(X\right)=\ln\left(B+1\right)$. However, for the sample space $x\in\mathbb{\mathbb{Z}}_{0}^{+}$, no such result exists because the entropy diverges to $\infty$ as $f_{X}\left(x\right)\rightarrow 0^{+}$ for all $x$, but the limiting values $f_{X}\left(x\right)=0$ do not form a proper probability distribution. Therefore, one must introduce a further constraint, such as $\textrm{E}_{X}\left[X^{\kappa}\right]\equiv\mu_{\kappa}$, subject to which entropy can be maximized. As noted above, if $\textrm{E}_{X}\left[X\right]\equiv\mu$, then the maximum-entropy PMF is that of the $\textrm{Geometric}\left(\varrho\right)$ distribution, $f_{X\mid\varrho}\left(x\right)=\varrho\left(1-\varrho\right)^{x}$, where $\varrho=\tfrac{1}{\mu+1}$ and $\textrm{H}_{F_{X\mid\varrho}}\left(X\right)=-\tfrac{1-\varrho}{\varrho}\ln\left(1-\varrho\right)-\ln\left(\varrho\right)$.
For a severity $Y\sim F_{Y}\left(y\right)$ defined on $y\in\mathbb{R}_{0}^{+}$, the Shannon entropy becomes
$\textrm{H}_{F_{Y}}\left(Y\right)=\textrm{E}_{F_{Y}}\left[-\ln\left(f_{Y}\left(Y\right)\right)\right]$
[…]
which often is called the differential entropy. Unfortunately, this quantity does not share all the properties of the expression in (7). Most notably, (8) can be negative, thereby failing to comport with the notion of expected “surprise” from a random outcome. Nevertheless, it still is possible to compare levels (higher vs. lower) of differential entropy among probability distributions, and to speak of maximum-entropy distributions.
If the sample space of $Y$ were bounded (e.g., $y\in\left[0,B\right)$), then, in a manner analogous to the discrete case, the associated entropy would be maximized for the $\textrm{Uniform}\left(0,B\right)$ PDF with $f_{Y\mid B}\left(y\right)=\tfrac{1}{B}$ for all $y$ and $\textrm{H}_{F_{Y\mid B}}\left(Y\right)=\ln\left(B\right)$. For $y\in\mathbb{R}_{0}^{+}$, one again must introduce a further constraint (e.g., $\textrm{E}_{Y}\left[Y^{\kappa}\right]\equiv\mu_{\kappa}$), as in the discrete case. As mentioned before, if $\textrm{E}_{Y}\left[Y\right]\equiv\mu$, then the maximum-entropy PDF is that of the $\textrm{Exponential}\left(\lambda\right)$ distribution, $f_{Y\mid\lambda}\left(y\right)=\lambda e^{-\lambda y}$, where $\lambda=\tfrac{1}{\mu}$ and $\textrm{H}_{F_{Y\mid\lambda}}\left(Y\right)=1-\ln\left(\lambda\right)$. For both the $\textrm{Uniform}\left(0,B\right)$ and $\textrm{Exponential}\left(\lambda\right)$ distributions, the differential entropy is negative for certain parameter values: $B\in\left(0,1\right)$ in the former case, and $\lambda\in\left(e,\infty\right)$ in the latter.
Given that we seek a nonnegative measure of functional simplicity, $\mathcal{S}\left(F_{0}\right)$, the Shannon entropy cannot be used directly. However, exponentiating this quantity yields a nonnegative alternative: the exponentiated entropy,
$e^{\textrm{H}_{F_{0}}\left(W\right)}=\exp\left(\textrm{E}_{F_{0}}\left[-\ln\left(f_{0}\left(W\right)\right)\right]\right),$ (9)
which preserves the order of $\textrm{H}_{F_{0}}\left(W\right)$ across different distribution functions. Moreover, this new quantity, like the Shannon entropy itself, does not depend directly on the number of parameters ($k$) characterizing $F_{0}$.
Rewriting (9) as
$\exp\left(\textrm{E}_{F_{0}}\left[\ln\left(\dfrac{1}{f_{0}\left(W\right)}\right)\right]\right)=\underset{p\rightarrow 0^{+}}{\lim}\left(\textrm{E}_{F_{0}}\left[\left|\dfrac{1}{f_{0}\left(W\right)}\right|^{p}\right]\right)^{1/p}$
reveals that the exponentiated entropy is essentially the geometric mean (or $0^{\textrm{th}}$ power mean) of $\tfrac{1}{f_{0}\left(W\right)}$ taken over $F_{0}\left(w\right)$. This is noteworthy because a closely related quantity, $\left(\textrm{E}_{F_{0}}\left[f_{0}\left(W\right)\right]\right)^{-1}$, the inverse of the Herfindahl-Hirschman index [^11] The Herfindahl-Hirschman index often is used as a measure of market concentration in regulatory economics. Although usually defined for discrete probability distributions, its extension to the continuous case is straightforward. (or $1^{\textrm{st}}$ power mean of $f_{0}\left(W\right)$), constitutes a natural measure of the dispersion of the probability distribution (see, e.g., Powers and Powers, 2015), for which greater values indicate a simpler, flatter PMF/PDF.
Recognizing that the $p^{\textrm{th}}$ power mean of $\tfrac{1}{f_{0}\left(W\right)}$ reasonably quantifies the distribution’s simplicity for all $p\in\mathbb{R}_{0}^{+}$, we tentatively consider all
$\mathcal{S}\left(F_{0}\right)=\left(\textrm{E}_{F_{0}}\left[\left|\dfrac{1}{f_{0}\left(W\right)}\right|^{p}\right]\right)^{1/p}$ (10)
as potential measures of functional simplicity.
4 Measuring Adaptability
4.1 Fisher Information
For the moment, assume $F_{0}\equiv F_{W\mid a}\left(w\right)$, where $a\in\mathbb{R}^{+}$ is a scalar parameter. To evaluate the functional adaptability of this probability distribution, it is natural to consider the impact of $a$ on the corresponding PMF or PDF, $f_{0}\left(w\right)$ – an effect that is measured by the Fisher information associated with a single observation of $W$:
$\textrm{I}_{F_{0}}\left(a\right)=\textrm{E}_{F_{0}}\left[\left(\dfrac{\partial\ln\left(f_{0}\left(W\right)\right)}{\partial a}\right)^{2}\right]$ (11)
$=-\textrm{E}_{F_{0}}\left[\dfrac{\partial^{2}\ln\left(f_{0}\left(W\right)\right)}{\partial a^{2}}\right].$
At a given value of $a$, the above equations quantify the extent by which changes in the parameter modify the shape of the PMF or PDF, on average. [^12] See Lehmann and Casella (1998) for a detailed exposition of Fisher information.
Like Shannon entropy, Fisher information is additive for independent random variables. [^13] Although Fisher information and Shannon entropy (information) possess certain similarities, they are designed to measure different types of “information” in different contexts, and therefore are not directly comparable. Nevertheless, they do share a mathematical connection through the calculation of relative entropy (Kullback-Leibler divergence). See, e.g., Cover and Thomas (2006). In particular, for a sample of $n$ independent and identically distributed (IID) observations, $\boldsymbol{W}=\left[W_{1},\ldots,W_{n}\right]$, (11) aggregates to
$\textrm{I}_{F_{0}}\left(a,n\right)={\displaystyle\sum_{i=1}^{n}}\textrm{E}_{F_{0}}\left[\left(\dfrac{\partial\ln\left(f_{0}\left(W_{i}\right)\right)}{\partial a}\right)^{2}\right]$
$=n\textrm{E}_{F_{0}}\left[\left(\dfrac{\partial\ln\left(f_{0}\left(W_{i}\right)\right)}{\partial a}\right)^{2}\right]$ (12)
$=-n\textrm{E}_{F_{0}}\left[\dfrac{\partial^{2}\ln\left(f_{0}\left(W_{i}\right)\right)}{\partial a^{2}}\right],$ (13)
and can be viewed as the amount of knowledge one can learn about the parameter $a$ by observing a random sample of size $n$. Although mathematically equivalent, the expressions in (12) and (13) offer different perspectives on this quantity.
In (12), the square of the first derivative captures the degree to which marginal changes in the parameter affect the shape of the PMF/PDF (at a fixed value of $a$), computed as an average over all possible values of $W$. In this case, a greater impact on the PMF/PDF, and thus the underlying probability distribution, implies that observations of $W$ tend to provide greater information about how the parameter characterizes the distribution. Since the PMF/PDF is transformed to the logarithmic scale, information grows in direct proportion to the sample size.
The expression in (13) reflects how steep, or pointed, the PMF/PDF is with respect to marginal changes in the parameter (at a fixed value of $a$) computed as an average over all possible values of $W$. Here, greater steepness (which is consistent with a greater impact of the parameter on the PMF/PDF) shows that the benefit of greater information about the parameter is accompanied by a notable shortcoming: small errors in estimating the true value of the parameter can lead to substantial errors in inference about the underlying probability distribution.
[…]
$\textrm{Var}{}_{F_{0}}\left[\hat{a}_{n}\right]\geq\dfrac{1}{\textrm{I}_{F_{0}}\left(a,n\right)}.$
Since increasing the sample size, $n$, results in a proportional increase in the Fisher information, it follows that the variance of any consistent estimator of $a$ will tend to decrease over $n$, while bounded below by the inverse of the Fisher information. Thus, in terms of changes in the sample size, greater information is associated with a smaller variance, which makes intuitive sense.
However, for a fixed sample size, one also can think of the relationship between the variance of the estimator and the Fisher information as permitting a tradeoff: for different parameterizations of the PMF/PDF, a parameterization that reduces the variance will imply an increase in the Fisher information. In this case, the increase in information is not beneficial because it means that any error in the estimate of $a$ that does occur will magnify certain errors in estimation of the relevant PMF/PDF.
To illustrate, consider an $\textrm{Exponential}\left(\lambda\right)$ random variable $Y$ with parameter $\lambda=6$, for which the asymptotic variance of the sample parameter, $\hat{\lambda}_{n}=\tfrac{1}{\bar{Y}_{n}}$, is $\tfrac{\lambda^{2}}{n}=\tfrac{36}{n}$. If we change the parameterization of $Y$ by writing the PDF in terms of the mean, $\theta=\tfrac{1}{\lambda}=\tfrac{1}{6}$, then the asymptotic variance of the sample estimator, $\hat{\theta}_{n}=\bar{Y}_{n}$, becomes much smaller: $\tfrac{\theta^{2}}{n}=\tfrac{1}{36n}$. However, using the latter estimator also causes many common calculations based on the estimate of the PDF, such as
$\Pr\left\{Y>1\right\}=e^{-\lambda}=e^{-1/\theta},$
to become much more sensitive to errors in the parameter estimator. For example, assuming an error of $+0.10$ in $\hat{\lambda}_{n}$, we would estimate the probability that $Y$ exceeds 1 as $0.0022$, which is fairly close to the true value of approximately $0.0025$. However, assuming the same error of $+0.10$ in $\hat{\theta}_{n}$, yields a corresponding probability estimate of $0.0235$, which is substantially farther from the true value.
4.2 $\boldsymbol{k^{\textrm{th}}}$ Root of Unnormalized Jeffreys Prior
For probability distributions with $k\geq 2$ parameters (i.e., $\boldsymbol{a}=\left[a_{1},\ldots,a_{k}\right]$), the Fisher information for $n=1$ generalizes to the $k\times k$ Fisher information matrix,
$\mathbf{I}_{F_{0}}^{\left(k\right)}\left(\boldsymbol{a}\right)=\textrm{E}_{F_{0}}\left[\begin{array}[]{ccc}\left(\dfrac{\partial\ln\left(f_{0}\left(W\right)\right)}{\partial a_{1}}\right)^{2}&\cdots&\left(\dfrac{\partial\ln\left(f_{0}\left(W\right)\right)}{\partial a_{1}}\right)\left(\dfrac{\partial\ln\left(f_{0}\left(W\right)\right)}{\partial a_{k}}\right)\\ \vdots&\ddots&\vdots\\ \left(\dfrac{\partial\ln\left(f_{0}\left(W\right)\right)}{\partial a_{k}}\right)\left(\dfrac{\partial\ln\left(f_{0}\left(W\right)\right)}{\partial a_{1}}\right)&\cdots&\left(\dfrac{\partial\ln\left(f_{0}\left(W\right)\right)}{\partial a_{k}}\right)^{2}\end{array}\right].$ (14)
Since the determinant of this matrix represents the volume spanned by the parallelepiped formed by the $k$ columns (or rows), it constitutes a natural extension to $k$ dimensions of the scalar measure of the average impact of $a$ on the PMF/PDF in (11). One then can raise this quantity to the $\tfrac{1}{2k}$ power to: (i) reduce its dimensional units from $k$ to 1, allowing it to be expressed on a “per-parameter” basis; and (ii) return the squared units formed by all matrix elements in (14) to their original scale (in the same way exponentiating the Shannon entropy did in Section 3). Interestingly, the resulting quantity, $\left(\det\left(\mathbf{I}_{F_{0}}^{\left(k\right)}\left(\boldsymbol{a}\right)\right)\right)^{1/2k}$, is equivalent to the $k^{\textrm{th}}$ root of the unnormalized Jeffreys prior PMF/PDF for the parameter vector $\boldsymbol{a}=\left[a_{1},\ldots,a_{k}\right]$:
$p_{\textrm{J}}\left(a_{1},\ldots,a_{k}\right)\propto\sqrt{\det\left(\mathbf{I}_{F_{0}}^{\left(k\right)}\left(\boldsymbol{a}\right)\right)}.$ (15)
Although $\left(\det\left(\mathbf{I}_{F_{0}}^{\left(k\right)}\left(\boldsymbol{a}\right)\right)\right)^{1/2k}$ appears to be a good candidate for measuring functional adaptability, a closer examination reveals a significant shortcoming. By rewriting (14) as
$\mathbf{I}_{F_{0}}^{\left(k\right)}\left(\boldsymbol{a}\right)=\textrm{E}_{F_{0}}\left[\left(\dfrac{1}{f_{0}\left(W\right)}\right)^{2}\mathbf{D}^{\left(k\right)}\left(\boldsymbol{a}\right)\right],$ (16)
with
$\mathbf{D}^{\left(k\right)}\left(\boldsymbol{a}\right)=\left[\begin{array}[]{ccc}\left(\dfrac{\partial f_{0}\left(W\right)}{\partial a_{1}}\right)^{2}&\cdots&\left(\dfrac{\partial f_{0}\left(W\right)}{\partial a_{1}}\right)\left(\dfrac{\partial f_{0}\left(W\right)}{\partial a_{k}}\right)\\ \vdots&\ddots&\vdots\\ \left(\dfrac{\partial f_{0}\left(W\right)}{\partial a_{k}}\right)\left(\dfrac{\partial f_{0}\left(W\right)}{\partial a_{1}}\right)&\cdots&\left(\dfrac{\partial f_{0}\left(W\right)}{\partial a_{k}}\right)^{2}\end{array}\right],$
one can see that
$\left(\det\left(\mathbf{I}_{F_{0}}^{\left(k\right)}\left(\boldsymbol{a}\right)\right)\right)^{1/2k}=\left(\det\left(\textrm{E}_{F_{0}}\left[\left(\dfrac{1}{f_{0}\left(W\right)}\right)^{2}\mathbf{D}^{\left(k\right)}\left(\boldsymbol{a}\right)\right]\right)\right)^{1/2k},$ (17)
where the component $\left(\tfrac{1}{f_{0}\left(W\right)}\right)^{2}$ is identical to that inside the expectation operation in (10) for $p=2$. Thus, (17) reflects not only the distribution’s adaptability (through $\mathbf{D}^{\left(k\right)}\left(\boldsymbol{a}\right)$), but also its simplicity, rendering it inappropriate for a measure of functional adaptability alone.
Rather fortuitously, this particular drawback actually identifies (17) as a good candidate for measuring the functional versatility of $F_{0}$, since such a measure must capture both simplicity and adaptability at the same time. Of course, the indicated expression is not unique in this regard. For example, one easily could modify (17) by replacing $\left(\tfrac{1}{f_{0}\left(W\right)}\right)^{2}$ with $\left|\tfrac{1}{f_{0}\left(W\right)}\right|^{p}$ for any $p\in\mathbb{R}_{0}^{+}$ and then raising the resulting determinant to the $\tfrac{1}{pk}$ power. However, (17) clearly offers the benefit of maximum analytical familiarity because it is based on the extensively studied Fisher information matrix.
Given that $\mathbf{I}_{F_{0}}^{\left(k\right)}\left(\boldsymbol{a}\right)$ depends on the parameter vector, $\boldsymbol{a}$, we still need to average over the prior distribution of $\boldsymbol{a}$ to ensure that the family of probability distributions specified by the null hypothesis in (1) is unconditionally versatile (i.e., across all values of $\boldsymbol{a}$). This yields the tentative measure
$\mathcal{V}\left(F_{0}\right)=\left(\det\left(\textrm{E}_{P}\left[\mathbf{I}_{F_{0}}^{\left(k\right)}\left(\boldsymbol{a}\right)\right]\right)\right)^{1/2k}$
$=\left(\det\left(\textrm{E}_{P}\left[\textrm{E}_{F_{0}}\left[\left(\dfrac{1}{f_{0}\left(W\right)}\right)^{2}\mathbf{D}^{\left(k\right)}\left(\boldsymbol{a}\right)\right]\right]\right)\right)^{1/2k},$ (18)
where $P\equiv P\left(\boldsymbol{a}\right)$ denotes the prior distribution function. The new component, $\textrm{E}_{P}\left[\mathbf{I}_{F_{0}}^{\left(k\right)}\left(\boldsymbol{a}\right)\right]$, is commonly called the Bayesian Fisher information matrix in the relevant literature. [^15] See, e.g., Daniels and Hogan (2008).
In the next section, we refine and formalize the versatility measure of (18). Then, in Section 6, we illustrate its usefulness in risk analysis through application to a number of common frequency and severity models.
5 Measuring Versatility
5.1 Maximum-Entropy Prior
The measure of functional versatility stated in (18) remains incomplete because its indicated prior distribution, $P\equiv P\left(\boldsymbol{a}\right)$, is unspecified. Here, as in many Bayesian settings with little or no information about distribution parameters, the choice of an appropriate prior distribution is somewhat challenging.
In the present context, all we know about the parameter vector $\boldsymbol{a}$ is that it should be representative of all possible vectors employed to characterize either a frequency or severity distribution. Since most frequency PMFs and severity PDFs are commonly formulated (or easily reformulated) with real-valued parameters spanning the space $\mathbb{R}^{+}$ (as in the Exponential case, where both the mean, $\theta$, and its inverse, $\lambda$, can be any positive real number), we will restrict attention to vectors $\boldsymbol{a}$ spanning $\left(\mathbb{R}^{+}\right)^{k}$ and seek a reasonable uninformative prior distribution according to the principle of indifference. This usually takes one of two approaches: the method of Jeffreys priors, or that of maximum-entropy priors. [^16] See Jaynes (2003) for an extensive discussion of uninformative prior distributions.
As observed in Subsection 4.2, the Jeffreys prior distribution is constructed so the joint PMF/PDF is proportional to the determinant of the square-root of the Fisher information matrix, as shown in (15). This formulation gives the prior distribution the desirable property that it is invariant with regard to an invertible and differentiable transformation of the parameter vector; in other words, if
$\left[b_{1},\ldots,b_{k}\right]=\boldsymbol{\textrm{g}}\left(a_{1},\ldots,a_{k}\right)$
for some invertible and differentiable $k$-dimensional vector function $\boldsymbol{\textrm{g}}\left(\boldsymbol{\cdot}\right)$, then
$p_{\textrm{J}}^{\prime}\left(b_{1},\ldots,b_{k}\right)=p_{\textrm{J}}\left(a_{1},\ldots,a_{k}\right)\left|\det\left(\boldsymbol{\textrm{J}}\right)\right|,$
where
$\boldsymbol{\textrm{J}}\equiv\left[\boldsymbol{\textrm{J}}_{i,j}\right]=\left[\dfrac{\partial a_{i}}{\partial b_{j}}\right]$
denotes the $k\times k$ Jacobian matrix of the inverse transformation.
[…]
The maximum-entropy approach, mentioned in Subsections 3.1 and 3.2, addresses the problem of no prior information by spreading the probability density as evenly as possible over a given sample space, subject to certain constraints. This technique always provides a proper PMF/PDF, but often involves some element of subjectivity in the selection of hyper-parameters (i.e., the parameters of the maximum-entropy distribution itself), either as fixed constants or random quantities with their own hyper-prior distributions. For simplicity, we will assume that it is reasonable to treat all hyper-parameters as fixed constants in the present analysis.
Among maximum-entropy distributions defined on $a\in\mathbb{R}^{+}$, the most obvious possibility is the Exponential distribution, which, as noted in Section 3, possesses the maximum-entropy PDF subject to a fixed mean, $\mu$. Unfortunately, the selection of the constant hyper-parameter $\mu$ poses serious difficulties. This is because the salient choice – $\mu=1$, based on the intuition that positive parameters can be divided into two mirror realms: real numbers greater than 1 and their inverses – actually divides $\mathbb{R}^{+}$ unequally, with
$\Pr\left\{a\leq 1\right\}=1-e^{-1/\mu}$
$=1-e^{-1}$
$\approx 0.6321.$
Furthermore, all other conspicuous choices for $\mu$ (such as $\mu=\tfrac{1}{\ln\left(2\right)}\approx 1.4427$, which implies a median of 1) appear rather unnatural.
To overcome this problem, we can require the distribution of $\ln\left(a\right)\in\mathbb{R}$ to be symmetric about 0. Here, the maximum-entropy distribution, subject to a fixed variance, $\sigma^{2}$, is $\textrm{Lognormal}\left(\mu=0,\sigma\right)$, with $\sigma\in\mathbb{R}^{+}$, and we can set the hyper-parameter $\sigma$ equal to 1 based on the rationale that $\ln\left(\sigma\right)$, like $\ln\left(a\right)$, must be symmetric about the origin (which, for a constant, is exactly 0). Taking the parameters $a_{1},\ldots,a_{k}$ in (18) to be IID $\textrm{Lognormal}\left(\mu=0,\sigma=1\right)$ random variables with joint distribution function $P_{\textrm{LN}}\equiv P_{\textrm{LN}}\left(\boldsymbol{a}\right)$ then yields the following expression:
$\mathcal{V}\left(F_{0}\right)=\left(\det\left(\textrm{E}_{\textrm{LN}}\left[\mathbf{I}_{F_{0}}^{\left(k\right)}\left(\boldsymbol{a}\right)\right]\right)\right)^{1/2k}.$ (19)
As a preliminary assessment of (19) for measuring distributional versatility, we apply the proposed measure to three 1-parameter distributions – $\textrm{Exponential}\left(\lambda\right)$, $\textrm{Gamma}\left(r,\lambda=1\right)$, and $\textrm{Weibull}\left(\lambda=1,\tau\right)$ – each of which constitutes a special case of the 3-parameter Generalized $\textrm{Generalized Gamma}\left(r,\lambda,\tau\right)$ (GG) family with PDF
$f_{Y\mid r,\lambda,\tau}^{\left(\Gamma\right)}\left(y\right)=\dfrac{\tau\lambda^{r}}{\Gamma\left(r\right)}y^{\tau r-1}e^{-\lambda y^{\tau}},$
for $r\in\mathbb{R}^{+}$, $\lambda\in\mathbb{R}^{+}$, and $\tau\in\mathbb{R}^{+}$. For each distribution, the third and fourth columns of Table 2 provide the mathematical expression for Fisher information and the computed value of $\mathcal{V}\left(F_{0}\right)$, respectively.
Table 2. Versatility Calculations for Three 1-Parameter Continuous Distributions
Distribution PDF Fisher $\boldsymbol{\mathcal{V}\left(F_{0}\right)}$
($\boldsymbol{F_{0}}$) Information
$\textrm{Exponential}\left(\lambda\right)$ $f_{Y\mid\lambda}^{\left(\textrm{E}\right)}\left(y\right)=\lambda e^{-\lambda y}$ $\dfrac{1}{\lambda^{2}}$ $2.7183$
$\equiv\textrm{GG}\left(r=1,\lambda,\tau=1\right)$
$\textrm{Gamma}\left(r,\lambda=1\right)$ $f_{Y\mid r,\lambda=1}^{\left(\Gamma\right)}\left(y\right)=\dfrac{y^{r-1}e^{-y}}{\Gamma\left(r\right)}$ $\psi^{\left(1\right)}\left(r\right)$ $2.8399$
$\equiv\textrm{GG}\left(r,\lambda=1,\tau=1\right)$
$\textrm{Weibull}\left(\lambda=1,\tau\right)$ $f_{Y\mid\lambda=1,\tau}^{\left(\textrm{Wei}\right)}\left(y\right)=\tau y^{\tau-1}e^{-y^{\tau}}$ $\dfrac{1}{\tau^{2}}\left[\left(1-\gamma\right)^{2}+\dfrac{\pi^{2}}{6}\right]$ $3.6709$
$\equiv\textrm{GG}\left(r=1,\lambda=1,\tau\right)$
Notes: (i) $\gamma=0.5772156649\ldots$ denotes Euler’s constant; and (ii) $\psi^{\left(1\right)}\left(r\right)=\tfrac{d^{2}\ln\left(\Gamma\left(r\right)\right)}{dr^{2}}$ denotes the polygamma function of order 1.
Examining the indicated values of $\mathcal{V}\left(F_{0}\right)$ in conjunction with representative plots of the corresponding PDFs in Figure 1 reveals a distinct pattern. Although all three distributions are relatively flat for most values of $y\in\mathbb{R}_{0}^{+}$, the versatility measure increases as the distribution is better able to concentrate probability around its mode.
For $\textrm{Exponential}\left(\lambda\right)$, the mode is always 0 and $f_{Y\mid\lambda}^{\left(\textrm{E}\right)}\left(0\right)=\lambda$ is finite. However, if the modes of $\textrm{Gamma}\left(r,\lambda=1\right)$ and $\textrm{Weibull}\left(\lambda=1,\tau\right)$ are 0 (which occurs for $r\leq 1$ and $\tau\leq 1$, respectively), then $\underset{y\rightarrow 0^{+}}{\lim}f_{Y\mid r,\lambda=1}^{\left(\Gamma\right)}\left(y\right)=\underset{y\rightarrow 0^{+}}{\lim}f_{Y\mid\lambda=1,\tau}^{\left(\textrm{Wei}\right)}\left(y\right)=\infty$ (except for the boundary cases of $r=1$ and $\tau=1$). Although $\textrm{Gamma}\left(r,\lambda=1\right)$ enjoys the flexibility of possessing an arbitrarily large mode (given by $r-1$ for $r\geq 1$), the associated value of $f_{Y\mid r,\lambda=1}^{\left(\Gamma\right)}\left(r-1\right)$ is bounded, decreasing from 1 (when $r=1$) to 0 (as $r\rightarrow\infty$). On the other hand, the mode of $\textrm{Weibull}\left(\lambda=1,\tau\right)$ (given by $\left(\tfrac{\tau-1}{\tau}\right)^{1/\tau}$ for $\tau\geq 1$), which is limited to the interval $\left[0,1\right)$, yields unbounded values of $f_{Y\mid\lambda=1,\tau}^{\left(\textrm{W}\right)}\left(\left(\tfrac{\tau-1}{\tau}\right)^{1/\tau}\right)$, increasing from 1 (when $\tau=1$) to $\infty$ (as $\tau\rightarrow\infty$).
[Uncaptioned image][Uncaptioned image][Uncaptioned image]
Figure 1. PDFs of $\textrm{Exponential}\left(\lambda\right)$, $\textrm{Gamma}\left(r,\lambda=1\right)$, and $\textrm{Weibull}\left(\lambda=1,\tau\right)$ Distributions
(parameter $=0.5$ for dotted line, $1.0$ for solid line, and $2.0$ for dashed line)
Thus, in moving from $\textrm{Exponential}\left(\lambda\right)$ to $\textrm{Gamma}\left(r,\lambda=1\right)$, the versatility measure increases because the enhanced adaptability from (i) a higher concentration of probability at the origin when $r<1$, and (ii) arbitrary placement of the mode when $r>1$, offsets the attendant decrease in simplicity from diminished entropy. Likewise, in going from $\textrm{Gamma}\left(r,\lambda=1\right)$ to $\textrm{Weibull}\left(\lambda=1,\tau\right)$, versatility continues to rise because the greater adaptability from a higher concentration of probability near $y=1$ (as $\tau\rightarrow\infty$), while retaining a similar concentration of probability at the origin (when $\tau<1$), offsets the further decrease in simplicity/entropy.
5.2 Least-Compressible Parameterization
Interestingly, the apparent reasonableness of the $\mathcal{V}\left(F_{0}\right)$ values in Table 2 is easily undermined by reparameterizing the indicated PDFs, subject to the restriction that each parameter span $\mathbb{R}^{+}$. For example, consider the two alternative parameterizations of the Exponential distribution:
$f_{Y\mid\lambda}^{\left(\textrm{E}\right)}\left(y\right)=\lambda e^{-\lambda y},$ (20)
as in Table 2; and
$f_{Y\mid\beta}^{\left(\textrm{E}\right)}\left(y\right)=\beta^{p}\exp\left(-\beta^{p}y\right),$ (21)
where we have replaced $\lambda$ by $\beta^{p}$ for some fixed value of $p\in\mathbb{R}^{+}$. As a result of this seemingly minor change, the value of $\mathcal{V}\left(F_{0}\right)$ is multiplied by a factor of $p$, which means that the versatility measure can be made either arbitrarily large (as $p\rightarrow\infty$) or arbitrarily small (as $p\rightarrow 0^{+}$).
Such sensitivity to a change in parameters renders the versatility measure of (19) useless unless we can find a way to invalidate variations like those in (21) for $p\neq 1$. Fortunately, this task is readily accomplished, at least conceptually, by comparing (21) with the more conventional parameterization of (20). Since (21) is clearly more complex (in the Kolmogorov sense) than (20) for any $p\neq 1$, we can remedy the problem of multiple parameterizations by requiring that only the least-compressible alternative (under some reasonable coding system for mathematical notation) be used to characterize a distribution function for measuring its versatility.
As mentioned in the discussion of notational complexity in Subsection 3.1, numerous methods for assessing the complexity of functional forms have been proposed in the literature, but none is broadly accepted or highly tractable. However, for the present purpose (which is much more delimited than the quest for a simplicity measure in the earlier subsection), it is possible to construct a relatively elementary – albeit informal – system for evaluating the compressibility of PMF/PDF parameterizations for commonly used frequency/severity distributions.
To this end, we will evaluate PMFs/PDFs according to the following rules for counting mathematical symbols (and exclude any functions incompatible with these rules):
$\bullet$ Equations must be written from left to right using the conventional order of common operations:
(1) operations within parentheses;
(2) exponentiation ($\wedge$) and taking roots ($\vee$);
(3) multiplication ($\times$) and division (///); and
(4) addition ($+$) and subtraction (--).
$\bullet$ Each individual real number (positive or negative), variable, and parameter counts as 1 symbol.
$\bullet$ Each common mathematical operation counts as 1 symbol.
$\bullet$ Each parenthesis counts as 1 symbol.
$\bullet$ Each elementary function (including the exponential, standard trigonometric, and hyperbolic functions, along with their respective inverses) counts as 1 symbol, written as $\phi$ (which denotes the given function’s unique symbol). However, the argument of any function must be placed in parentheses, which count as 2 symbols.
$\bullet$ The gamma function counts as 1 symbol, written as $\Gamma$ [^18] Although generally considered a non-elementary function, the gamma function is explicitly included in our system because of its frequent use in probability modeling. The beta function, which also is commonly used, is excluded because it can be written as a simple combination of gamma functions.
$\bullet$ Other special functions, as well as more complicated mathematical operations (e.g., summations, limits, derivatives, integrals, etc.) are not permitted.
Applying the above system to the PDFs in (20) and (21) yields the following results:
$f_{Y\mid\lambda}^{\left(\textrm{E}\right)}\left(y\right)=\lambda\times e\left(\tilde{1}\times\lambda\times y\right)$ (22)
(where $e\left(\cdot\right)$ denotes the exponential function and $\tilde{1}$ denotes $-1$), which requires 10 symbols; and
$f_{Y\mid\beta}^{\left(\textrm{E}\right)}\left(y\right)=\left(\beta\wedge p\right)\times e\left(\tilde{1}\times\beta\wedge p\times y\right),$ (23)
which requires 16 symbols. Thus, according to the proposed evaluation method, the parameterization given by (20) clearly is more compressible than that of (21), compelling us to use the former for computing $\mathcal{V}\left(F_{0}\right)$ (unless we can find an even less-compressible alternative).
Naturally, it is possible to write a given PMF/PDF so that it requires more symbols than necessary. For example, one could re-express (22) as
$f_{Y\mid\lambda}^{\left(\textrm{E}\right)}\left(y\right)=\lambda\times\left(\left(e\wedge\left(\tilde{1}\right)\right)\wedge\lambda\right)\wedge y$ (24)
and (23) as
$f_{Y\mid\beta}^{\left(\textrm{E}\right)}\left(y\right)=\left(\beta\wedge p\right)\times\left(\left(e\wedge\left(\tilde{1}\right)\right)\wedge\left(\beta\wedge p\right)\right)\wedge y,$ (25)
requiring 15 and 23 symbols, respectively. To avoid this problem, we assume that all PMFs/PDFs are written as efficiently as possible.
Furthermore, it is possible for two or more distinct parameterizations to minimize the symbol count under the system outlined above. For example, setting $\lambda=\tfrac{1}{\theta}$ gives the PDF
$f_{Y\mid\theta}^{\left(\textrm{E}\right)}\left(y\right)=\dfrac{1}{\theta}e^{-y/\theta},$
which was used in Subsection 3.1. Rewriting this PDF as
$f_{Y\mid\theta}^{\left(\textrm{E}\right)}\left(y\right)=e\left(\tilde{1}\times y/\theta\right)/\theta$
reveals that it requires 10 symbols, the same number as the PDF of (20). Although the two corresponding values of $\mathcal{V}\left(F_{0}\right)$ happen to be equal in this case, they sometimes may be different. If that occurs, then the easiest way to resolve the ambiguity is to average the indicated versatility measures. (See, e.g., the Negative $\textrm{Negative Binomial}\left(r,m\right)$ and Discrete $\textrm{Discrete Weibull}\left(m,\tau\right)$ examples of Subsection 6.2 below.)
Given the preceding discussion of this subsection, we now modify the functional versatility measure of (19) as follows:
$\mathcal{V}\left(F_{0}\right)=\left(\det\left(\textrm{E}_{\textrm{LN}}\left[\mathbf{I}_{F_{0}}^{\left(k\right)}\left(\boldsymbol{a^{*}}\right)\right]\right)\right)^{1/2k},$ (26)
where $\boldsymbol{a^{*}}$ denotes the parameter vector associated with the least-compressible parameterization of $F_{W\mid\boldsymbol{a}}^{\left(0\right)}\left(w\right)$.
6 Illustrative Applications
Having defined a formal measure of functional versatility, we now apply this measure to eight distinct families of 2-parameter distributions to explore its implications. Of these eight families, four are defined on $\mathbb{R}_{0}^{+}$ (i.e., appropriate for severities) and four on $\mathbb{Z}_{0}^{+}$ (i.e., appropriate for frequencies).
6.1 Continuous Distributions
The four continuous families – $\textrm{Gamma}\left(r,\lambda\right)$, $\textrm{Weibull}\left(\lambda,\tau\right)$, Pareto $\textrm{Pareto 2}\left(\alpha,\vartheta\right)$, and $\textrm{Lognormal}\left(\ln\left(\nu\right),\sigma\right)$ – are selected primarily for their common use as severity models. Their PDFs and versatility measures are presented in Table 3, with all parameters taken to span $\mathbb{R}^{+}$. To provide some (admittedly limited) sense of the impact of individual parameters on the versatility measure, the table includes two 1-parameter special cases of each 2-parameter family (three of which are identical to the 1-parameter examples shown in Table 2 of Subsection 5.1).
A brief review of Table 3 yields the following observations:
  • The versatility measures associated with the Pareto $\textrm{Pareto 2}\left(\alpha,\vartheta\right)$ family are among the lowest in the table. Given that these PDFs, like that of the $\textrm{Exponential}\left(\lambda\right)$ distribution, are relatively flat and strictly decreasing, the low versatility measures suggest that the simplicity of the PDFs is offset by their limited adaptability.
  • For three of the four families, the versatility measure of the 2-parameter distribution falls between the measures of the two 1-parameter special cases. This property is somewhat intuitive because one would expect the distributional characteristics contributed by each of the two parameters individually to be “averaged” when taken together. Therefore, it is important to note that the property does not hold for all distributions. For $\textrm{Gamma}\left(r,\lambda\right)$, the fact that
    $\mathcal{V}\left(F_{Y\mid r,\lambda}^{\left(\Gamma\right)}\right)>\max\left\{\mathcal{V}\left(F_{Y\mid r=1,\lambda}^{\left(\Gamma\right)}\right),\mathcal{V}\left(F_{Y\mid r,\lambda=1}^{\left(\Gamma\right)}\right)\right\}$
    suggests that the $r$ and $\lambda$ parameters positively reinforce each other’s impact on versatility. However, the analysis of such marginal contributions to versatility lies beyond the scope of the present work.
  • The parameters with a direct effect on tail behavior – that is, $\tau$ for the $\textrm{Weibull}\left(\lambda,\tau\right)$ distribution and $\alpha$ for the Pareto $\textrm{Pareto 2}\left(\alpha,\vartheta\right)$ distribution – appear to have relatively larger effects on versatility than other parameters.
  • The versatility measure of $\textrm{Lognormal}\left(\ln\left(\nu\right),\sigma=1\right)$ is considerably lower than other measures in the table. This is because the PDF is quite flat for $\nu>1$, with the function’s simplicity offset by its limited adaptability, and quite concentrated (in the neighborhood of 0) for $\nu<1$, with the function’s adaptability offset by its inadequate simplicity. Although we do not propose a specific minimum value of the versatility measure for model selection, it appears that $\mathcal{V}\left(F_{Y\mid\ln\left(\nu\right),\sigma=1}^{\left(\textrm{LN}\right)}\right)=1.0000$ is a strong indicator of idiosyncrasy in this case.
Table 3. Versatility Calculations for Four 2-Parameter Continuous Distributions and Special Cases
Distribution PDF $\boldsymbol{\mathcal{V}\left(F_{0}\right)}$
($\boldsymbol{F_{0}}$)
(1) $\left(1\right)\textrm{ Gamma}\left(r,\lambda\right)$ $f_{Y\mid r,\lambda}^{\left(\Gamma\right)}\left(y\right)=\dfrac{\lambda^{r}y^{r-1}e^{-\lambda y}}{\Gamma\left(r\right)}$ $3.1264$
$\textrm{Gamma}\left(r=1,\lambda\right)\equiv\textrm{Exponential}\left(\lambda\right)$ $f_{Y\mid r=1,\lambda}^{\left(\Gamma\right)}\left(y\right)=\lambda e^{-\lambda y}$ $2.7183$
$\textrm{Gamma}\left(r,\lambda=1\right)$ $f_{Y\mid r,\lambda=1}^{\left(\Gamma\right)}\left(y\right)=\dfrac{y^{r-1}e^{-y}}{\Gamma\left(r\right)}$ $2.8399$
(2) $\left(2\right)\textrm{ Weibull}\left(\lambda,\tau\right)$ $f_{Y\mid\lambda,\tau}^{\left(\textrm{Wei}\right)}\left(y\right)=\tau\lambda y^{\tau-1}e^{-\lambda y^{\tau}}$ $3.4349$
$\textrm{Weibull}\left(\lambda,\tau=1\right)\equiv\textrm{Exponential}\left(\lambda\right)$ $f_{Y\mid\lambda,\tau=1}^{\left(\textrm{Wei}\right)}\left(y\right)=\lambda e^{-\lambda y}$ $2.7183$
$\textrm{Weibull}\left(\lambda=1,\tau\right)$ $f_{Y\mid\lambda=1,\tau}^{\left(\textrm{Wei}\right)}\left(y\right)=\tau y^{\tau-1}e^{-y^{\tau}}$ $3.6709$
(3) Pareto $\left(3\right)\textrm{ Pareto 2}\left(\alpha,\vartheta\right)$ $f_{Y\mid\alpha,\vartheta}^{\left(\textrm{P2}\right)}\left(y\right)=\dfrac{\alpha\vartheta^{\alpha}}{\left(y+\vartheta\right)^{\alpha+1}}$ $2.0874$
Pareto $\textrm{Pareto 2}\left(\alpha=1,\vartheta\right)$ $f_{Y\mid\alpha=1,\vartheta}^{\left(\textrm{P2}\right)}\left(y\right)=\dfrac{\vartheta}{\left(y+\vartheta\right)^{2}}$ $1.5694$
Pareto $\textrm{Pareto 2}\left(\alpha,\vartheta=1\right)$ $f_{Y\mid\alpha,\vartheta=1}^{\left(\textrm{P2}\right)}\left(y\right)=\dfrac{\alpha}{\left(y+1\right)^{\alpha+1}}$ $2.7183$
(4) $\left(4\right)\textrm{ Lognormal}\left(\ln\left(\nu\right),\sigma\right)$ $f_{Y\mid\ln\left(\nu\right),\sigma}^{\left(\textrm{LN}\right)}\left(y\right)=\dfrac{\exp\left(-\tfrac{\left(\ln\left(y\right)-\ln\left(\nu\right)\right)^{2}}{2\sigma^{2}}\right)}{\sqrt{2\pi}\sigma y}$ $3.2327$
$\textrm{Lognormal}\left(\ln\left(\nu\right)=0,\sigma\right)$ $f_{Y\mid\ln\left(\nu\right)=0,\sigma}^{\left(\textrm{LN}\right)}\left(y\right)=\dfrac{\exp\left(-\tfrac{\left(\ln\left(y\right)\right)^{2}}{2\sigma^{2}}\right)}{\sqrt{2\pi}\sigma y}$ $3.8440$
$\textrm{Lognormal}\left(\ln\left(\nu\right),\sigma=1\right)$ $f_{Y\mid\ln\left(\nu\right),\sigma=1}^{\left(\textrm{LN}\right)}\left(y\right)=\dfrac{\exp\left(-\tfrac{\left(\ln\left(y\right)-\ln\left(\nu\right)\right)^{2}}{2}\right)}{\sqrt{2\pi}y}$ $1.0000$
6.2 Discrete Distributions
Table 4 presents the PMFs and versatility measures of four discrete families – Negative $\textrm{Negative Binomial}\left(r,\tfrac{m}{m+1}\right)$, Discrete $\textrm{Discrete Weibull}\left(\tfrac{m}{m+1},\tau\right)$, $\textrm{Waring}\left(\alpha,\vartheta\right)$, and Generalized $\textrm{Generalized Poisson}\left(\lambda,\varsigma\right)$ – with all parameters spanning $\mathbb{R}^{+}$. The first three of these families serve as respective analogues of the first three continuous families discussed above. In particular:
(1) Negative $\textrm{Negative Binomial}\left(r,\tfrac{m}{m+1}\right)$ models the sum of IID inter-arrival times from a Bernoulli process in the same manner as $\textrm{Gamma}\left(r,\lambda\right)$ models the sum of IID inter-arrival times from a Poisson process;
(2) Discrete $\textrm{Discrete Weibull}\left(\tfrac{m}{m+1},\tau\right)$ [^19] See Nakagawa and Osaki (1975). possesses a tail parameter, $\tau$, with a functional role similar to that of the corresponding $\textrm{Weibull}\left(\lambda,\tau\right)$ parameter, and also shares the memoryless $\textrm{Geometric}\left(\tfrac{m}{m+1}\right)$ distribution as a special case with Negative $\textrm{Negative Binomial}\left(r,\tfrac{m}{m+1}\right)$ in the same way $\textrm{Weibull}\left(\lambda,\tau\right)$ shares the memoryless $\textrm{Exponential}\left(\lambda\right)$ distribution with $\textrm{Gamma}\left(r,\lambda\right)$; and
(3) $\textrm{Waring}\left(\alpha,\vartheta\right)$ is characterized by strictly decreasing PMFs with heavy tails (such that for all real $\kappa\leq\alpha$, $\textrm{E}_{X\mid\alpha,\vartheta}\left[X^{\kappa}\right]=\infty$), just as Pareto $\textrm{Pareto 2}\left(\alpha,\vartheta\right)$ is characterized by strictly decreasing PDFs with heavy tails (such that for all real $\kappa\leq\alpha$, $\textrm{E}_{Y\mid\alpha,\vartheta}\left[Y^{\kappa}\right]=\infty$).
Although the fourth discrete family, Generalized $\textrm{Generalized Poisson}\left(\lambda,\varsigma\right)$, is not directly analogous to $\textrm{Lognormal}\left(\ln\left(\nu\right),\sigma\right)$, both of these families often serve as well-known benchmarks for the study of other 2-parameter models.
A comparison of Tables 3 and 4 reveals:
  • For the Negative $\textrm{Negative Binomial}\left(r,\tfrac{m}{m+1}\right)$ family with $\textrm{E}_{X\mid r,\tfrac{m}{m+1}}\left[X\right]=\tfrac{1}{m}$, there is a second, equally compressible parameterization, Negative $\textrm{Negative Binomial}\left(r,\tfrac{1}{m+1}\right)$, for which $\textrm{E}_{X\mid r,\tfrac{1}{m+1}}\left[X\right]=m$. This is analogous to the fact that $\textrm{Gamma}\left(r,\lambda\right)$ possesses the equally compressible counterpart $\textrm{Gamma}\left(r,\theta\right)$ in the continuous case (as noted in Subsection 5.2). However, unlike the continuous case, the two different parameterizations yield different versatility measures, which we resolve by averaging (as suggested in Subsection 5.2). The same issue arises for the Discrete $\textrm{Discrete Weibull}\left(\tfrac{m}{m+1},\tau\right)$ family, and is handled in the same way. [^20] The versatility measures of distributions parameterized by $\tfrac{1}{m+1}$ are noticeably smaller than those of distributions based on $\tfrac{m}{m+1}$ because the former PMFs are particularly sensitive to $m$ for values of $x$ close to 00, whereas the latter are relatively more sensitive as $x\rightarrow\infty$. For the discrete sample space $x\in\mathbb{Z}_{0}^{+}$, there is little opportunity for a PMF to benefit from functional adaptability in the neighborhood of 00, hence less functional versatility.
Table 4. Versatility Calculations for Four 2-Parameter Discrete Distributions and Special Cases
Distribution PMF $\boldsymbol{\mathcal{V}\left(F_{0}\right)}$
($\boldsymbol{F_{0}}$)
[…]
Average $1.4816$
[…]
Average $1.7850$
[…]
Average $2.0585$
[…]
Average $1.7850$
[…]
(3) $\left(3\right)\textrm{ Waring}\left(\alpha,\vartheta\right)$ $f_{X\mid\alpha,\vartheta}^{\left(\textrm{War}\right)}\left(x\right)=\dfrac{\alpha\Gamma\left(\alpha+\vartheta\right)\Gamma\left(x+\vartheta\right)}{\Gamma\left(\vartheta\right)\Gamma\left(x+\alpha+\vartheta+1\right)}$ $1.3997$
$\textrm{Waring}\left(\alpha=1,\vartheta\right)$ $f_{X\mid\alpha=1,\vartheta}^{\left(\textrm{War}\right)}\left(x\right)=\dfrac{\vartheta}{\left(x+\vartheta\right)\left(x+\vartheta+1\right)}$ $0.9423$
$\textrm{Waring}\left(\alpha,\vartheta=1\right)$ $f_{X\mid\alpha,\vartheta=1}^{\left(\textrm{War}\right)}\left(x\right)=\dfrac{\alpha\Gamma\left(\alpha+1\right)\Gamma\left(x+1\right)}{\Gamma\left(x+\alpha+2\right)}$ $2.2441$
[…]
  • The versatility measures associated with the first three discrete families are consistently lower than the corresponding measures of their continuous analogues. This arises from the same phenomenon mentioned in Footnote 20: the discrete sample space $x\in\mathbb{Z}_{0}^{+}$ affords dramatically less opportunity for functional adaptability in the neighborhood of 00 than does the continuous sample space $y\in\mathbb{R}_{0}^{+}$.
  • For all four families, the versatility measure of the 2-parameter distribution falls between the measures of the two 1-parameter special cases.
  • The parameters with a direct effect on tail behavior – that is, $\tau$ for the averaged Discrete $\textrm{Weibull}\left(\tfrac{m}{m+1},\tau\right)$ and Discrete $\textrm{Discrete Weibull}\left(\tfrac{1}{m+1},\tau\right)$ distributions and $\alpha$ for the $\textrm{Waring}\left(\alpha,\vartheta\right)$ distribution – appear to have relatively larger effects on versatility than other parameters.
  • The versatility measures of Negative $\textrm{Negative Binomial}\left(r,\tfrac{m}{m+1}=\tfrac{1}{2}\right)$ and $\textrm{Waring}\left(\alpha=1,\vartheta\right)$ are considerably lower than other measures in the table. This is because the two PMFs are quite flat for all values of $r$ and $\vartheta$, respectively, with both functions’ simplicity offset by limited adaptability. As noted in conjunction with $\textrm{Lognormal}\left(\ln\left(\nu\right),\sigma=1\right)$ in the previous subsection, we do not propose a specific minimum cutoff for the versatility measure; however, it appears that both $\mathcal{V}\left(F_{X\mid r,\tfrac{m}{m+1}=\tfrac{1}{2}}^{\left(\textrm{NB}\right)}\right)=1.0151$ and $\mathcal{V}\left(F_{X\mid\alpha=1,\vartheta}^{\left(\textrm{War}\right)}\right)=0.9423$ are strong indicators of idiosyncrasy.
7 Conclusion
In the present study, we have proposed an approach to reduce the idiosyncratic overfitting of risk-analytic models by ensuring that, for a fixed number of parameters, the probability distributions used to model frequencies and severities are reasonably versatile. After developing the concept of versatility based on underlying components of functional simplicity and adaptability, we derived a mathematical measure of this property (in the form of a normalized Bayesian Fisher information matrix) in (19). This measure was refined to account for parameterization differences by requiring, in (26), that the relevant PMF or PDF be expressed in its least-compressible form.
The three simple (1-parameter) severity models of Table 2 revealed that, for relatively flat distributions, the versatility measure increases as the distribution is better able to concentrate probability around its mode. Subsequently, we explored the implications of this measure for a variety of 2-parameter severity and frequency distributions in Tables 3 and 4, respectively. These examples showed that:
  • The versatility measure of a 2-parameter distribution often falls between the measures of two 1-parameter special cases (with each of the parameters fixed in turn), suggesting that the characteristics contributed by the two parameters individually are “averaged” when taken together.
  • Parameters with a direct effect on tail behavior – for example, $\tau$ for the $\textrm{Weibull}\left(\lambda,\tau\right)$ and Discrete $\textrm{Weibull}\left(\tfrac{m}{m+1},\tau\right)$ distributions and $\alpha$ for the Pareto $\textrm{Pareto 2}\left(\alpha,\vartheta\right)$ and $\textrm{Waring}\left(\alpha,\vartheta\right)$ distributions – appear to have relatively larger impacts on versatility than other parameters.
  • The versatility measures associated with certain discrete families tend to be lower than those of their continuous analogues because versatility is sensitive to the level of functional adaptability afforded in the neighborhood of 00.
  • The versatility measures of certain distributions (i.e., $\textrm{Lognormal}\left(\ln\left(\nu\right),\sigma=1\right)$, Negative $\textrm{Binomial}\left(r,\tfrac{m}{m+1}=\tfrac{1}{2}\right)$, and $\textrm{Waring}\left(\alpha=1,\vartheta\right)$) are considerably lower than other measures, raising serious concerns about idiosyncratic behavior.
Overall, we would argue that the proposed versatility measure may be employed to assess the potential usefulness of probability distributions a priori by accounting for and mitigating problems of idiosyncratic overfitting. Although our analysis does not propose minimum cutoff values for model-selection purposes, the versatility measure can be used as an effective guide for identifying a range of acceptable values based on the measures arising from distribution families found to be useful in practice. For example, the relatively low measures associated with the Pareto $\textrm{Pareto 2}\left(\alpha,\vartheta\right)$ family should not necessarily disqualify this commonly used model, but rather suggest a lower region for reasonable versatility levels. On the other hand, the substantially lower measures computed for the $\textrm{Lognormal}\left(\ln\left(\nu\right),\sigma=1\right)$, Negative $\textrm{Negative Binomial}\left(r,\tfrac{m}{m+1}=\tfrac{1}{2}\right)$, and $\textrm{Waring}\left(\alpha=1,\vartheta\right)$ distributions likely suggest these models should be avoided.
Much further work remains to be done in this area, particularly involving the analysis of marginal contributions to versatility of additional parameters for $k\geq 2$. In this regard, it seems likely that the most useful potential applications of the proposed versatility measure lie in the investigation and comparison of probability distributions within multiple-parameter families; for example, the Generalized Beta, Burr, Pareto, and Pearson families increasingly used to model severities in risk analysis.

### v3

@@
 $H_{0}:\>W\sim F_{0}\equiv F_{W\mid\boldsymbol{a}}^{\left(0\right)}\left(w\right),$ (1)
-where $F_{0}\equiv F_{W\mid\boldsymbol{\alpha}}^{\left(0\right)}\left(w\right)$ represents the cumulative distribution function of either $W=X\in\mathbb{\mathbb{Z}}_{0}^{+}$ or $W=Y\in\mathbb{R}_{0}^{+}$, with unknown $k$-dimensional parameter vector $\boldsymbol{a}=\left[a_{1},\ldots,a_{k}\right]$.
+where $F_{0}\equiv F_{W\mid\boldsymbol{a}}^{\left(0\right)}\left(w\right)$ represents the cumulative distribution function of either $W=X\in\mathbb{\mathbb{Z}}_{0}^{+}$ or $W=Y\in\mathbb{R}_{0}^{+}$, with unknown $k$-dimensional parameter vector $\boldsymbol{a}=\left[a_{1},\ldots,a_{k}\right]$.
 The following subsections address characteristics of $F_{0}$ that are particularly desirable in this setting.
@@
 To clarify these relationships, let the real-valued quantities $\mathcal{I}\left(F_{0}\right)\geq 0$, $\mathcal{C}\left(F_{0}\right)\geq 0,$ and $\mathcal{T}\left(F_{0}\right)\geq 0$ denote hypothetical heuristic measures of the intrinsic idiosyncrasy, complexity, and contrivance of the distribution function $F_{0}$, where each quantity is assumed to be independent of the number of parameters, $k$.
-In addition, set $\mathcal{V}\left(F_{0}\right)=1/\mathcal{I}\left(F_{0}\right),$ $\mathcal{S}\left(F_{0}\right)=1/\mathcal{C}\left(F_{0}\right),$ and $\mathcal{A}\left(F_{0}\right)=1/\mathcal{T}\left(F_{0}\right)$ as the corresponding inverse (i.e., opposite) quantities: versatility, simplicity, and adaptability, respectively.
-Then, from the discussion of the preceding paragraph, it follows (heuristically) that
-$\mathcal{I}\left(F_{0}\right)=\mathcal{I}_{1}\left(\mathcal{C}^{\left(+\right)},\mathcal{T}^{\left(+\right)}\right)$
-$=\mathcal{I}_{2}\left(\mathcal{S}^{\left(-\right)},\mathcal{A}^{\left(-\right)}\right),$ (2)
-where the superscripts indicate the direction of the argument’s relationship with $\mathcal{I}$ (i.e., $\left(+\right)=\textrm{positive}$ or $\left(-\right)=\textrm{negative}$).
+In addition, set $\mathcal{V}\left(F_{0}\right)=\tfrac{1}{\mathcal{I}\left(F_{0}\right)},$ $\mathcal{S}\left(F_{0}\right)=\tfrac{1}{\mathcal{C}\left(F_{0}\right)},$ and $\mathcal{A}\left(F_{0}\right)=\tfrac{1}{\mathcal{T}\left(F_{0}\right)}$ as the corresponding inverse (i.e., opposite) quantities: versatility, simplicity, and adaptability, respectively.
+Then, from the discussion of the preceding paragraph, it follows that $\mathcal{I}\left(F_{0}\right)$ may be viewed (heuristically) as a bivariate function of $\mathcal{C}\left(F_{0}\right)$ and $\mathcal{T}\left(F_{0}\right)$ (or alternatively, of $\mathcal{S}\left(F_{0}\right)$ and $\mathcal{A}\left(F_{0}\right)$); that is,
+$\mathcal{I}\left(F_{0}\right)=\mathcal{I}\left(\mathcal{C}^{\left(+\right)},\mathcal{T}^{\left(+\right)}\right)$
+$=\mathcal{I}\left(\mathcal{S}^{\left(-\right)},\mathcal{A}^{\left(-\right)}\right),$ (2)
+where the superscripts indicate the direction of the argument’s impact on $\mathcal{I}$ (i.e., $\left(+\right)=\textrm{positive}$ and $\left(-\right)=\textrm{negative}$).
 Now let $\mathcal{P}\left(F_{0}\right)\geq 0$ denote a hypothetical real-valued measure of the parsimony of $F_{0}$, taken to be negatively related to both its number of parameters and functional complexity (because both quantities oppose the objective of simplicity), but independent of functional contrivance (because unnaturally manipulated functions can possess simple or complex structures).
-We then can write (heuristically)
-$\mathcal{P}\left(F_{0}\right)=\mathcal{P}_{1}\left(k^{\left(-\right)},\mathcal{C}^{\left(-\right)}\right)$
-$=\mathcal{P}_{2}\left(k^{\left(-\right)},\mathcal{S}^{\left(+\right)}\right).$ (3)
+We then can treat $\mathcal{P}\left(F_{0}\right)$ (heuristically) as a bivariate function of $k$ and $\mathcal{C}\left(F_{0}\right)$ (or alternatively, of $k$ and $\mathcal{S}\left(F_{0}\right)$):
+$\mathcal{P}\left(F_{0}\right)=\mathcal{P}\left(k^{\left(-\right)},\mathcal{C}^{\left(-\right)}\right)$
+$=\mathcal{P}\left(k^{\left(-\right)},\mathcal{S}^{\left(+\right)}\right){\color[rgb]{1,0,0},}$ (3)
+where the superscripts again indicate the direction of the argument’s impact.
 Functional simplicity, $\mathcal{S}\left(F_{0}\right)$, which both increases parsimony and reduces idiosyncrasy, will be investigated further in Section 3.
@@
 Letting $\mathcal{R}\left(F_{0}\right)\geq 0$ denote a hypothetical real-valued measure of the robustness of $F_{0}$, we will take this quantity to be positively related to its number of parameters (more of which permit greater accuracy of fit), negatively related to its degree of functional contrivance (because unnaturally manipulated functions are less adaptable to a wide range of empirical observations), and independent of functional complexity (because unnaturally complicated functions can enjoy various degrees of adaptability).
-We thus write (again heuristically)
-$\mathcal{R}\left(F_{0}\right)=\mathcal{R}_{1}\left(k^{\left(+\right)},\mathcal{T}^{\left(-\right)}\right)$
-$=\mathcal{R}_{2}\left(k^{\left(+\right)},\mathcal{A}^{\left(+\right)}\right).$ (4)
+We thus can view $\mathcal{R}\left(F_{0}\right)$ (heuristically) as a bivariate function of $k$ and $\mathcal{T}\left(F_{0}\right)$ (or alternatively, of $k$ and $\mathcal{A}\left(F_{0}\right)$):
+$\mathcal{R}\left(F_{0}\right)=\mathcal{R}\left(k^{\left(+\right)},\mathcal{T}^{\left(-\right)}\right)$
+$=\mathcal{R}\left(k^{\left(+\right)},\mathcal{A}^{\left(+\right)}\right).$ (4)
 Functional adaptability, $\mathcal{A}\left(F_{0}\right)$, which both increases robustness and reduces idiosyncrasy, will be studied in Section 4.
@@
 However, as indicated by equations (3) and (4), these quantities may depend explicitly on $k$, in which case they can be increased simply by adjusting the number of parameters, rather than by increasing functional simplicity and adaptability (the quantities directly relevant to idiosyncrasy).
-With this in mind, we treat functional versatility as the inverse of idiosyncrasy:
-$\mathcal{V}\left(F_{0}\right)=\mathcal{V}\left(\mathcal{S}^{\left(+\right)},\mathcal{A}^{\left(+\right)}\right),$ (5)
+With this in mind, we will focus on functional versatility as the inverse of idiosyncrasy:
+$\mathcal{V}\left(F_{0}\right)=\dfrac{1}{\mathcal{I}\left(\mathcal{S}^{\left(-\right)},\mathcal{A}^{\left(-\right)}\right)}$
+$=\mathcal{V}\left(\mathcal{S}^{\left(+\right)},\mathcal{A}^{\left(+\right)}\right),$ (5)
 and propose a formal expression for this quantity in Section 5.
@@
 Concept Heuristic Measure Inverse
-Simplicity $\mathcal{S}\left(F_{0}\right)=1/\mathcal{C}\left(F_{0}\right)$ Complexity
-Parsimony $\mathcal{P}\left(F_{0}\right)=\mathcal{P}_{2}\left(\mathcal{S}^{\left(+\right)},k^{\left(-\right)}\right)$
-$=\mathcal{P}_{1}\left(\mathcal{C}^{\left(-\right)},k^{\left(-\right)}\right)$
-Adaptability $\mathcal{A}\left(F_{0}\right)=1/\mathcal{T}\left(F_{0}\right)$ Contrivance
-Robustness $\mathcal{R}\left(F_{0}\right)=\mathcal{R}_{2}\left(\mathcal{A}^{\left(+\right)},k^{\left(+\right)}\right)$
-$=\mathcal{R}_{1}\left(\mathcal{T}^{\left(-\right)},k^{\left(+\right)}\right)$
+Simplicity $\mathcal{S}\left(F_{0}\right)=\dfrac{1}{\mathcal{C}\left(F_{0}\right)}$ Complexity
+Parsimony $\mathcal{P}\left(F_{0}\right)=\mathcal{P}\left(k^{\left(-\right)},\mathcal{S}^{\left(+\right)}\right)$
+$=\mathcal{P}\left(k^{\left(-\right)},\mathcal{C}^{\left(-\right)}\right)$
+Adaptability $\mathcal{A}\left(F_{0}\right)=\dfrac{1}{\mathcal{T}\left(F_{0}\right)}$ Contrivance
+Robustness $\mathcal{R}\left(F_{0}\right)=\mathcal{R}\left(k^{\left(+\right)},\mathcal{A}^{\left(+\right)}\right)$
+$=\mathcal{R}\left(k^{\left(+\right)},\mathcal{T}^{\left(-\right)}\right)$
 Versatility $\mathcal{V}\left(F_{0}\right)=\mathcal{V}\left(\mathcal{S}^{\left(+\right)},\mathcal{A}^{\left(+\right)}\right)$ Idiosyncrasy
-$=1/\mathcal{I}\left(F_{0}\right)$
+$=\dfrac{1}{\mathcal{I}\left(F_{0}\right)}$
 3 Measuring Simplicity
@@
 Certainly, functions can be devised to achieve high levels of Kolmogorov (algorithmic) complexity (also called Kolmogorov entropy), meaning that they cannot be compressed into short string lengths in a mathematical coding system.
-[^8] See Li and Vitányi (2019) for a recent treatment of Kolmogorov complexity/entropy.
+[^8] See Li and Vitányi (2019) for a general overview of Kolmogorov complexity/entropy and Lu and Oliveira (2022) for insights into current applications.
 However, such constructions necessitate intentionally injecting elements of disorder or disorganization, since functions formed according to an ordered plan or pattern (even if somewhat complicated) possess less Kolmogorov complexity because they are more compressible.
 In the case of Type B, we encounter an apparent paradox: PMFs/PDFs with greater Shannon (information-theoretic) entropy (also called Shannon information) tend to possess simpler, flatter shapes.
-[^9] See Gray (2011) for a comprehensive overview of Shannon entropy.
+[^9] See Gray (2011) for a broad exposition of Shannon entropy and Nanda and Chowdhury (2021) for a survey of major developments up to the present.
 For example, the maximum-entropy discrete distribution on $x\in\mathbb{\mathbb{Z}}_{0}^{+}$, subject to a fixed mean, is given by the $\textrm{Geometric}\left(\varrho\right)$ PMF ($f_{X\mid\varrho}\left(x\right)=\varrho\left(1-\varrho\right)^{x}$, for $\varrho\in\left(0,1\right)$), and the maximum-entropy continuous distribution on $y\in\mathbb{R}_{0}^{+}$, again subject to a fixed mean, is given by the $\textrm{Exponential}\left(\lambda\right)$ PDF ($f_{Y\mid\lambda}\left(y\right)=\lambda e^{-\lambda y}$, for $\lambda\in\mathbb{R}^{+}$).
@@
 4.1 Fisher Information
-For the moment, assume $F_{0}\equiv F_{W\mid a}\left(w\right)$, where $a\in\mathbb{R}^{+}$ is a scalar parameter.
+For the moment, assume $F_{0}\equiv F_{W\mid a}^{\left(0\right)}\left(w\right)$, where $a$ is a scalar parameter belonging to some open set $\mathbb{A}\subseteq\mathbb{R}$.
 To evaluate the functional adaptability of this probability distribution, it is natural to consider the impact of $a$ on the corresponding PMF or PDF, $f_{0}\left(w\right)$ – an effect that is measured by the Fisher information associated with a single observation of $W$:
-$\textrm{I}_{F_{0}}\left(a\right)=\textrm{E}_{F_{0}}\left[\left(\dfrac{\partial\ln\left(f_{0}\left(W\right)\right)}{\partial a}\right)^{2}\right]$ (11)
-$=-\textrm{E}_{F_{0}}\left[\dfrac{\partial^{2}\ln\left(f_{0}\left(W\right)\right)}{\partial a^{2}}\right].$
-At a given value of $a$, the above equations quantify the extent by which changes in the parameter modify the shape of the PMF or PDF, on average.
-[^12] See Lehmann and Casella (1998) for a detailed exposition of Fisher information.
+$\textrm{I}_{F_{0}}\left(a\right)=\textrm{E}_{F_{0}}\left[\left(\dfrac{\partial\ln\left(f_{0}\left(W\right)\right)}{\partial a}\right)^{2}\right]$
+$=-\textrm{E}_{F_{0}}\left[\dfrac{\partial^{2}\ln\left(f_{0}\left(W\right)\right)}{\partial a^{2}}\right],$ (11)
+[…]
 Like Shannon entropy, Fisher information is additive for independent random variables.
@@
 See, e.g., Cover and Thomas (2006).
-In particular, for a sample of $n$ independent and identically distributed (IID) observations, $\boldsymbol{W}=\left[W_{1},\ldots,W_{n}\right]$, (11) aggregates to
+In particular, for a sample of $n$ independent and identically distributed (IID) observations, $\boldsymbol{W}=\left[W_{1},W_{2},\ldots,W_{n}\right]$, (11) aggregates to
 $\textrm{I}_{F_{0}}\left(a,n\right)={\displaystyle\sum_{i=1}^{n}}\textrm{E}_{F_{0}}\left[\left(\dfrac{\partial\ln\left(f_{0}\left(W_{i}\right)\right)}{\partial a}\right)^{2}\right]$
@@
 Here, greater steepness (which is consistent with a greater impact of the parameter on the PMF/PDF) shows that the benefit of greater information about the parameter is accompanied by a notable shortcoming: small errors in estimating the true value of the parameter can lead to substantial errors in inference about the underlying probability distribution.
-[…]
+The interplay of negative and positive implications of the Fisher information becomes apparent in the context of the Cramér-Rao lower bound on the asymptotic variance of any consistent estimator, $\hat{a}_{n}$, of $a$: [^14] A consistent estimator, $\hat{a}_{n}$, converges in probability to the true parameter, $a$.
+This is denoted by $\hat{a}_{n}\overset{\mathcal{P}}{\rightarrow}a$; that is, for any $\epsilon>0$, $\underset{n\rightarrow\infty}{\lim}\Pr\left\{\left|\hat{a}_{n}-a\right|>\epsilon\right\}=0$.
 $\textrm{Var}{}_{F_{0}}\left[\hat{a}_{n}\right]\geq\dfrac{1}{\textrm{I}_{F_{0}}\left(a,n\right)}.$
@@
 to become much more sensitive to errors in the parameter estimator.
-For example, assuming an error of $+0.10$ in $\hat{\lambda}_{n}$, we would estimate the probability that $Y$ exceeds 1 as $0.0022$, which is fairly close to the true value of approximately $0.0025$.
-However, assuming the same error of $+0.10$ in $\hat{\theta}_{n}$, yields a corresponding probability estimate of $0.0235$, which is substantially farther from the true value.
+For example, assuming an error of $+0.1000$ in $\hat{\lambda}_{n}$, we would estimate $\Pr\left\{Y>1\right\}$ as $0.0022$, which falls short of the true value ($\Pr\left\{Y>1\right\}\approx 0.0025$) by $0.0003$.
+To achieve a comparably small deviation in the estimate of $\Pr\left\{Y>1\right\}$ under the alternative parameterization would require the absolute error in $\hat{\theta}_{n}$ be no greater than
+$\max\left\{\left|\dfrac{1}{-\ln\left(0.0025-0.0003\right)}-\dfrac{1}{6}\right|,\left|\dfrac{1}{-\ln\left(0.0025+0.0003\right)}-\dfrac{1}{6}\right|\right\}\approx\max\left\{\left|-0.0032\right|,\left|0.0035\right|\right\}$
+$=0.0035.$
+Computation of the Fisher information frequently arises in the context of maximum-likelihood estimation (MLE) because the estimator generated, $\hat{a}_{n}$, is (under the regularity conditions identified in Footnote 12) asymptotically Normal, [^15] An asymptotically Normal estimator, $\hat{a}_{n}$, converges in distribution to the Normal distribution with mean $a$ (the true parameter) and variance equal to the Cramér-Rao lower bound.
+This is denoted by $\hat{a}_{n}\overset{\mathcal{D}}{\rightarrow}\textrm{Normal}\left(a,\>\tfrac{1}{\textrm{I}_{F_{0}}\left(a,n\right)}\right)$; that is, $\underset{n\rightarrow\infty}{\lim}\left|\Pr\left\{\tfrac{\hat{a}_{n}-a}{\sqrt{1/\textrm{I}_{F_{0}}\left(a,n\right)}}\leq t\right\}-\Phi\left(t\right)\right|=0$ for all $t\in\mathbb{R}$, where $\Phi\left(\cdot\right)$ is the CDF of the Standard Normal distribution.
+permitting its variance to be estimated well by the inverse of the Fisher information; that is,
+$\textrm{Var}{}_{F_{0}}\left[\hat{a}_{n}\right]\approx\dfrac{1}{\textrm{I}_{F_{0}}\left(\hat{a}_{n},n\right)}$
+for large $n,$ where $\textrm{I}_{F_{0}}\left(\hat{a}_{n},n\right)$ denotes an estimator of $\textrm{I}_{F_{0}}\left(a,n\right)$ obtained by substituting $\hat{a}_{n}$ for $a$.
+Two common obstacles to employing the Fisher information in this way are: (1) the failure of $W\sim F_{0}\left(w\right)$ to satisfy the first condition in Footnote 12 because one or both endpoints of the sample space ($\mathbb{W}$) depend on $a$; and (2) the non-concavity of the log-likelihood, ${\textstyle\sum_{i=1}^{n}}\ln\left(f_{0}\left(W_{i}\right)\right)$, as a function of $a$.
 4.2 $\boldsymbol{k^{\textrm{th}}}$ Root of Unnormalized Jeffreys Prior
 For probability distributions with $k\geq 2$ parameters (i.e., $\boldsymbol{a}=\left[a_{1},\ldots,a_{k}\right]$), the Fisher information for $n=1$ generalizes to the $k\times k$ Fisher information matrix,
-$\mathbf{I}_{F_{0}}^{\left(k\right)}\left(\boldsymbol{a}\right)=\textrm{E}_{F_{0}}\left[\begin{array}[]{ccc}\left(\dfrac{\partial\ln\left(f_{0}\left(W\right)\right)}{\partial a_{1}}\right)^{2}&\cdots&\left(\dfrac{\partial\ln\left(f_{0}\left(W\right)\right)}{\partial a_{1}}\right)\left(\dfrac{\partial\ln\left(f_{0}\left(W\right)\right)}{\partial a_{k}}\right)\\ \vdots&\ddots&\vdots\\ \left(\dfrac{\partial\ln\left(f_{0}\left(W\right)\right)}{\partial a_{k}}\right)\left(\dfrac{\partial\ln\left(f_{0}\left(W\right)\right)}{\partial a_{1}}\right)&\cdots&\left(\dfrac{\partial\ln\left(f_{0}\left(W\right)\right)}{\partial a_{k}}\right)^{2}\end{array}\right].$ (14)
+$\mathbf{I}_{F_{0}}^{\left(k\right)}\left(\boldsymbol{a}\right)=\textrm{E}_{F_{0}}\left[\begin{array}[]{ccc}\left(\dfrac{\partial\ln\left(f_{0}\left(W\right)\right)}{\partial a_{1}}\right)^{2}&\cdots&\left(\dfrac{\partial\ln\left(f_{0}\left(W\right)\right)}{\partial a_{1}}\right)\left(\dfrac{\partial\ln\left(f_{0}\left(W\right)\right)}{\partial a_{k}}\right)\\ \vdots&\ddots&\vdots\\ \left(\dfrac{\partial\ln\left(f_{0}\left(W\right)\right)}{\partial a_{k}}\right)\left(\dfrac{\partial\ln\left(f_{0}\left(W\right)\right)}{\partial a_{1}}\right)&\cdots&\left(\dfrac{\partial\ln\left(f_{0}\left(W\right)\right)}{\partial a_{k}}\right)^{2}\end{array}\right]{\color[rgb]{1,0,0},}$ (14)
+which is well defined subject to natural extensions of the regularity conditions stated in Footnote 12.
 Since the determinant of this matrix represents the volume spanned by the parallelepiped formed by the $k$ columns (or rows), it constitutes a natural extension to $k$ dimensions of the scalar measure of the average impact of $a$ on the PMF/PDF in (11).
 One then can raise this quantity to the $\tfrac{1}{2k}$ power to: (i) reduce its dimensional units from $k$ to 1, allowing it to be expressed on a “per-parameter” basis; and (ii) return the squared units formed by all matrix elements in (14) to their original scale (in the same way exponentiating the Shannon entropy did in Section 3).
-Interestingly, the resulting quantity, $\left(\det\left(\mathbf{I}_{F_{0}}^{\left(k\right)}\left(\boldsymbol{a}\right)\right)\right)^{1/2k}$, is equivalent to the $k^{\textrm{th}}$ root of the unnormalized Jeffreys prior PMF/PDF for the parameter vector $\boldsymbol{a}=\left[a_{1},\ldots,a_{k}\right]$:
+Interestingly, the resulting quantity, $\left(\det\left(\mathbf{I}_{F_{0}}^{\left(k\right)}\left(\boldsymbol{a}\right)\right)\right)^{1/2k}$, is equivalent to the $k^{\textrm{th}}$ root of the unnormalized Jeffreys prior PMF/PDF for the parameter vector $\boldsymbol{a}=\left[a_{1},a_{2},\ldots,a_{k}\right]$:
 $p_{\textrm{J}}\left(a_{1},\ldots,a_{k}\right)\propto\sqrt{\det\left(\mathbf{I}_{F_{0}}^{\left(k\right)}\left(\boldsymbol{a}\right)\right)}.$ (15)
@@
 This usually takes one of two approaches: the method of Jeffreys priors, or that of maximum-entropy priors.
-[^16] See Jaynes (2003) for an extensive discussion of uninformative prior distributions.
+[^17] See Jaynes (2003) for a classic treatment of uninformative prior distributions and Llorente et al.
+(2023) for some contemporary perspectives.
 As observed in Subsection 4.2, the Jeffreys prior distribution is constructed so the joint PMF/PDF is proportional to the determinant of the square-root of the Fisher information matrix, as shown in (15).
 This formulation gives the prior distribution the desirable property that it is invariant with regard to an invertible and differentiable transformation of the parameter vector; in other words, if
-$\left[b_{1},\ldots,b_{k}\right]=\boldsymbol{\textrm{g}}\left(a_{1},\ldots,a_{k}\right)$
+$\left[b_{1},b_{2},\ldots,b_{k}\right]=\boldsymbol{\textrm{g}}\left(a_{1},a_{2},\ldots,a_{k}\right)$
 for some invertible and differentiable $k$-dimensional vector function $\boldsymbol{\textrm{g}}\left(\boldsymbol{\cdot}\right)$, then
-$p_{\textrm{J}}^{\prime}\left(b_{1},\ldots,b_{k}\right)=p_{\textrm{J}}\left(a_{1},\ldots,a_{k}\right)\left|\det\left(\boldsymbol{\textrm{J}}\right)\right|,$
+$p_{\textrm{J}}^{\prime}\left(b_{1},b_{2},\ldots,b_{k}\right)=p_{\textrm{J}}\left(a_{1},a_{2},\ldots,a_{k}\right)\left|\det\left(\boldsymbol{\textrm{J}}\right)\right|,$
 where
@@
 Here, the maximum-entropy distribution, subject to a fixed variance, $\sigma^{2}$, is $\textrm{Lognormal}\left(\mu=0,\sigma\right)$, with $\sigma\in\mathbb{R}^{+}$, and we can set the hyper-parameter $\sigma$ equal to 1 based on the rationale that $\ln\left(\sigma\right)$, like $\ln\left(a\right)$, must be symmetric about the origin (which, for a constant, is exactly 0).
-Taking the parameters $a_{1},\ldots,a_{k}$ in (18) to be IID $\textrm{Lognormal}\left(\mu=0,\sigma=1\right)$ random variables with joint distribution function $P_{\textrm{LN}}\equiv P_{\textrm{LN}}\left(\boldsymbol{a}\right)$ then yields the following expression:
+Taking the parameters $a_{1},a_{2},\ldots,a_{k}$ in (18) to be IID $\textrm{Lognormal}\left(\mu=0,\sigma=1\right)$ random variables with joint distribution function $P_{\textrm{LN}}\equiv P_{\textrm{LN}}\left(\boldsymbol{a}\right)$ then yields the following expression:
 $\mathcal{V}\left(F_{0}\right)=\left(\det\left(\textrm{E}_{\textrm{LN}}\left[\mathbf{I}_{F_{0}}^{\left(k\right)}\left(\boldsymbol{a}\right)\right]\right)\right)^{1/2k}.$ (19)
+Since $\textrm{E}_{\textrm{LN}}\left[\left|a_{1}^{p_{1}}a_{2}^{p_{2}}\cdots a_{k}^{p_{k}}\right|\right]<\infty$ for all $\boldsymbol{p}\in\mathbb{R}^{k}$, the right-hand side of (19) is well defined as long as the individual elements of $\mathbf{I}_{F_{0}}^{\left(k\right)}\left(\boldsymbol{a}\right)$ do not grow faster than $\left|a_{1}^{p_{1}}a_{2}^{p_{2}}\cdots a_{k}^{p_{k}}\right|$ as the individual $a_{i}$ approach 00 or $\infty$.
 As a preliminary assessment of (19) for measuring distributional versatility, we apply the proposed measure to three 1-parameter distributions – $\textrm{Exponential}\left(\lambda\right)$, $\textrm{Gamma}\left(r,\lambda=1\right)$, and $\textrm{Weibull}\left(\lambda=1,\tau\right)$ – each of which constitutes a special case of the 3-parameter Generalized $\textrm{Generalized Gamma}\left(r,\lambda,\tau\right)$ (GG) family with PDF
@@
 On the other hand, the mode of $\textrm{Weibull}\left(\lambda=1,\tau\right)$ (given by $\left(\tfrac{\tau-1}{\tau}\right)^{1/\tau}$ for $\tau\geq 1$), which is limited to the interval $\left[0,1\right)$, yields unbounded values of $f_{Y\mid\lambda=1,\tau}^{\left(\textrm{W}\right)}\left(\left(\tfrac{\tau-1}{\tau}\right)^{1/\tau}\right)$, increasing from 1 (when $\tau=1$) to $\infty$ (as $\tau\rightarrow\infty$).
-[Uncaptioned image][Uncaptioned image][Uncaptioned image]
-Figure 1. PDFs of $\textrm{Exponential}\left(\lambda\right)$, $\textrm{Gamma}\left(r,\lambda=1\right)$, and $\textrm{Weibull}\left(\lambda=1,\tau\right)$ Distributions
-(parameter $=0.5$ for dotted line, $1.0$ for solid line, and $2.0$ for dashed line)
 Thus, in moving from $\textrm{Exponential}\left(\lambda\right)$ to $\textrm{Gamma}\left(r,\lambda=1\right)$, the versatility measure increases because the enhanced adaptability from (i) a higher concentration of probability at the origin when $r<1$, and (ii) arbitrary placement of the mode when $r>1$, offsets the attendant decrease in simplicity from diminished entropy.
 Likewise, in going from $\textrm{Gamma}\left(r,\lambda=1\right)$ to $\textrm{Weibull}\left(\lambda=1,\tau\right)$, versatility continues to rise because the greater adaptability from a higher concentration of probability near $y=1$ (as $\tau\rightarrow\infty$), while retaining a similar concentration of probability at the origin (when $\tau<1$), offsets the further decrease in simplicity/entropy.
+[Uncaptioned image][Uncaptioned image][Uncaptioned image]
+Figure 1. PDFs of the $\textrm{Exponential}\left(\lambda\right)$, $\textrm{Gamma}\left(r,\lambda=1\right)$, and $\textrm{Weibull}\left(\lambda=1,\tau\right)$ Distributions (from left to right)
+[variable parameter $=0.5$ for dotted line, $1.0$ for solid line, and $2.0$ for dashed line]
+Given the definition in (19), it is clear that $\mathcal{V}\left(F_{0}\right)\in\mathbb{R}^{+},$ and thus not bounded above.
+This makes the interpretation of versatility magnitudes somewhat challenging, and raises the natural question of whether it is possible to transpose $\mathcal{V}\left(F_{0}\right)$ to a bounded interval, such as $\left(0,1\right)$, in a way that affords meaningful interpretations (as in the case of the regression $R^{2}$, which can be viewed as the proportion of variation explained by a posited regression model).
+One salient approach would be to consider transformations $\textrm{T}\left(\mathcal{V}\left(F_{0}\right)\right):\mathbb{R}^{+}\rightarrow\left(0,1\right)$ that endow a percentile to each possible versatility measure.
+In other words, if we could define and identify a theoretical distribution, $F_{\mathcal{V}}\left(v\right)$, of $\mathcal{V}\left(F_{0}\right)$ over all probability distributions $F_{0}$ within some prescribed category (say, e.g., severity distributions on $\mathbb{R}_{0}^{+}$), then we might set $\textrm{T}\left(\mathcal{V}\left(F_{0}\right)\right)=100F_{\mathcal{V}}^{-1}\left(\mathcal{V}\left(F_{0}\right)\right)$.
+Although the problem of identifying a suitable theoretical distribution $F_{\mathcal{V}}\left(v\right)$ appears largely intractable, there are potential alternatives.
+For example, one could employ an uninformative prior on $\mathbb{R}^{+}$ (such as $\textrm{Lognormal}\left(\mu=0,\sigma=1\right)$), or base the choice of $F_{\mathcal{V}}\left(v\right)$ on empirical, rather than theoretical, considerations.
+In fact, the latter approach is likely to arise naturally in applied work by observing a large number of versatility measures for a certain category of distributions $F_{0}$, yielding an empirical distribution of $\mathcal{V}\left(F_{0}\right)$.
+For purposes of the current study, we will defer a deeper investigation of this issue to future research, recognizing that actual computations of $\mathcal{V}\left(F_{0}\right)$ are useful primarily for comparing the characteristics of different probability distributions, rather than assessing the qualities of any single distribution in isolation.
+In this way, the proposed versatility measure is comparable to other unbounded assessment measures, such as Akaike’s Information Criterion and Mallows’ $C_{p}$ Statistic.
 5.2 Least-Compressible Parameterization
@@
 $\bullet$ Other special functions, as well as more complicated mathematical operations (e.g., summations, limits, derivatives, integrals, etc.) are not permitted.
+[^20] We note that this rule precludes applications to mixture distributions, which are commonly used in risk analysis.
+However, two practical considerations mitigate the impact of this restriction: (1) it remains feasible to compute and compare the versatility measures of different mixture distributions whose formulations in terms of integrals/sums over the mixing distribution are carried out in a consistent manner; and (2) it is possible to assess the versatility of a mixture distribution informally by measuring the versatilities of its individual components (i.e., the kernel and mixing distribution) separately.
 Applying the above system to the PDFs in (20) and (21) yields the following results:
@@
 Having defined a formal measure of functional versatility, we now apply this measure to eight distinct families of 2-parameter distributions to explore its implications.
-Of these eight families, four are defined on $\mathbb{R}_{0}^{+}$ (i.e., appropriate for severities) and four on $\mathbb{Z}_{0}^{+}$ (i.e., appropriate for frequencies).
+Four of these families are defined on $\mathbb{R}_{0}^{+}$, and thus appropriate for modeling severities, and the other four are defined on $\mathbb{Z}_{0}^{+}$, and appropriate for frequencies.
+Following this overview, we will investigate one of the continuous families more closely to show how the versatility measure can be employed in practice to select certain probability distributions over others.
 6.1 Continuous Distributions
-The four continuous families – $\textrm{Gamma}\left(r,\lambda\right)$, $\textrm{Weibull}\left(\lambda,\tau\right)$, Pareto $\textrm{Pareto 2}\left(\alpha,\vartheta\right)$, and $\textrm{Lognormal}\left(\ln\left(\nu\right),\sigma\right)$ – are selected primarily for their common use as severity models.
+The four continuous families – $\textrm{Gamma}\left(r,\lambda\right)$, $\textrm{Weibull}\left(\lambda,\tau\right)$, Pareto $\textrm{Pareto 2}\left(\alpha,\vartheta\right)$, and $\textrm{Lognormal}\left(\ln\left(\nu\right),\sigma\right)$ – are chosen primarily for their common use as severity models.
 Their PDFs and versatility measures are presented in Table 3, with all parameters taken to span $\mathbb{R}^{+}$.
@@
   • For three of the four families, the versatility measure of the 2-parameter distribution falls between the measures of the two 1-parameter special cases.
-This property is somewhat intuitive because one would expect the distributional characteristics contributed by each of the two parameters individually to be “averaged” when taken together.
-Therefore, it is important to note that the property does not hold for all distributions.
-For $\textrm{Gamma}\left(r,\lambda\right)$, the fact that
-    $\mathcal{V}\left(F_{Y\mid r,\lambda}^{\left(\Gamma\right)}\right)>\max\left\{\mathcal{V}\left(F_{Y\mid r=1,\lambda}^{\left(\Gamma\right)}\right),\mathcal{V}\left(F_{Y\mid r,\lambda=1}^{\left(\Gamma\right)}\right)\right\}$
-    suggests that the $r$ and $\lambda$ parameters positively reinforce each other’s impact on versatility.
-However, the analysis of such marginal contributions to versatility lies beyond the scope of the present work.
+This property seems intuitive because one would expect the distributional characteristics contributed by each of the two parameters individually to be “averaged” when taken together.
+However, it is important to note that this property does not always hold for extreme values of the “fixed” parameter in the 1-parameter distributions.
+For example, as the fixed parameter $\alpha^{*}\rightarrow\infty$ in the 1-parameter Pareto $\textrm{Pareto 2}\left(\alpha^{*},\vartheta\right)$, the versatility measure converges to $2.7183$ (which is greater than the measure associated with Pareto $\textrm{Pareto 2}\left(\alpha,\vartheta\right)$).
+Furthermore, as the fixed parameter $\sigma^{*}\rightarrow 0$ in the 1-parameter $\textrm{Lognormal}\left(\ln\left(\nu\right),\sigma^{*}\right)$, the versatility measure diverges to $\infty$ (well above the measure associated with $\textrm{Lognormal}\left(\ln\left(\nu\right),\sigma\right)$).
+Apart from these cautionary observations, the analysis of marginal contributions to versatility from individual parameters lies beyond the scope of the present work.
   • The parameters with a direct effect on tail behavior – that is, $\tau$ for the $\textrm{Weibull}\left(\lambda,\tau\right)$ distribution and $\alpha$ for the Pareto $\textrm{Pareto 2}\left(\alpha,\vartheta\right)$ distribution – appear to have relatively larger effects on versatility than other parameters.
@@
 possesses a tail parameter, $\tau$, with a functional role similar to that of the corresponding $\textrm{Weibull}\left(\lambda,\tau\right)$ parameter, and also shares the memoryless $\textrm{Geometric}\left(\tfrac{m}{m+1}\right)$ distribution as a special case with Negative $\textrm{Negative Binomial}\left(r,\tfrac{m}{m+1}\right)$ in the same way $\textrm{Weibull}\left(\lambda,\tau\right)$ shares the memoryless $\textrm{Exponential}\left(\lambda\right)$ distribution with $\textrm{Gamma}\left(r,\lambda\right)$; and
-(3) $\textrm{Waring}\left(\alpha,\vartheta\right)$ is characterized by strictly decreasing PMFs with heavy tails (such that for all real $\kappa\leq\alpha$, $\textrm{E}_{X\mid\alpha,\vartheta}\left[X^{\kappa}\right]=\infty$), just as Pareto $\textrm{Pareto 2}\left(\alpha,\vartheta\right)$ is characterized by strictly decreasing PDFs with heavy tails (such that for all real $\kappa\leq\alpha$, $\textrm{E}_{Y\mid\alpha,\vartheta}\left[Y^{\kappa}\right]=\infty$).
+(3) $\textrm{Waring}\left(\alpha,\vartheta\right)$ is characterized by strictly decreasing PMFs with heavy tails (such that for all real $\kappa\leq\alpha$, $\textrm{E}_{X\mid\alpha,\vartheta}^{\left(\textrm{War}\right)}\left[X^{\kappa}\right]=\infty$), just as Pareto $\textrm{Pareto 2}\left(\alpha,\vartheta\right)$ is characterized by strictly decreasing PDFs with heavy tails (such that for all real $\kappa\leq\alpha$, $\textrm{E}_{Y\mid\alpha,\vartheta}^{\left(\textrm{P2}\right)}\left[Y^{\kappa}\right]=\infty$).
 Although the fourth discrete family, Generalized $\textrm{Generalized Poisson}\left(\lambda,\varsigma\right)$, is not directly analogous to $\textrm{Lognormal}\left(\ln\left(\nu\right),\sigma\right)$, both of these families often serve as well-known benchmarks for the study of other 2-parameter models.
-A comparison of Tables 3 and 4 reveals:
+From Table 4, we make the following observations:
   • For the Negative $\textrm{Negative Binomial}\left(r,\tfrac{m}{m+1}\right)$ family with $\textrm{E}_{X\mid r,\tfrac{m}{m+1}}\left[X\right]=\tfrac{1}{m}$, there is a second, equally compressible parameterization, Negative $\textrm{Negative Binomial}\left(r,\tfrac{1}{m+1}\right)$, for which $\textrm{E}_{X\mid r,\tfrac{1}{m+1}}\left[X\right]=m$.
@@
 The same issue arises for the Discrete $\textrm{Discrete Weibull}\left(\tfrac{m}{m+1},\tau\right)$ family, and is handled in the same way.
+The versatility measures associated with the first three discrete families are consistently lower than the corresponding measures of their continuous analogues.
+This arises from the same phenomenon mentioned in Footnote 20: the discrete sample space $x\in\mathbb{Z}_{0}^{+}$ affords dramatically less opportunity for functional adaptability in the neighborhood of 00 than does the continuous sample space $y\in\mathbb{R}_{0}^{+}$.
+  • For all four families, the versatility measure of the 2-parameter distribution falls between the measures of the two 1-parameter special cases.
+However, as noted with regard to Table 3, this property does not always hold for extreme values of the fixed parameter in the 1-parameter distributions.
+  • The parameters with a direct effect on tail behavior – that is, $\tau$ for the averaged Discrete $\textrm{Weibull}\left(\tfrac{m}{m+1},\tau\right)$ and Discrete $\textrm{Discrete Weibull}\left(\tfrac{1}{m+1},\tau\right)$ distributions and $\alpha$ for the $\textrm{Waring}\left(\alpha,\vartheta\right)$ distribution – appear to have relatively larger effects on versatility than other parameters.
 [^22] The versatility measures of distributions parameterized by $\tfrac{1}{m+1}$ are noticeably smaller than those of distributions based on $\tfrac{m}{m+1}$ because the former PMFs are particularly sensitive to $m$ for values of $x$ close to 00, whereas the latter are relatively more sensitive as $x\rightarrow\infty$.
@@
 […]
-  • The versatility measures associated with the first three discrete families are consistently lower than the corresponding measures of their continuous analogues.
-This arises from the same phenomenon mentioned in Footnote 20: the discrete sample space $x\in\mathbb{Z}_{0}^{+}$ affords dramatically less opportunity for functional adaptability in the neighborhood of 00 than does the continuous sample space $y\in\mathbb{R}_{0}^{+}$.
-  • For all four families, the versatility measure of the 2-parameter distribution falls between the measures of the two 1-parameter special cases.
-  • The parameters with a direct effect on tail behavior – that is, $\tau$ for the averaged Discrete $\textrm{Weibull}\left(\tfrac{m}{m+1},\tau\right)$ and Discrete $\textrm{Discrete Weibull}\left(\tfrac{1}{m+1},\tau\right)$ distributions and $\alpha$ for the $\textrm{Waring}\left(\alpha,\vartheta\right)$ distribution – appear to have relatively larger effects on versatility than other parameters.
-  • The versatility measures of Negative $\textrm{Negative Binomial}\left(r,\tfrac{m}{m+1}=\tfrac{1}{2}\right)$ and $\textrm{Waring}\left(\alpha=1,\vartheta\right)$ are considerably lower than other measures in the table.
+  • The versatility measures of Negative $\textrm{Negative Binomial}\left(r,\tfrac{m}{m+1}=\tfrac{1}{2}\right)$ and $\textrm{Waring}\left(\alpha=1,\vartheta\right)$ are noticeably lower than other measures in the table.
 This is because the two PMFs are quite flat for all values of $r$ and $\vartheta$, respectively, with both functions’ simplicity offset by limited adaptability.
 As noted in conjunction with $\textrm{Lognormal}\left(\ln\left(\nu\right),\sigma=1\right)$ in the previous subsection, we do not propose a specific minimum cutoff for the versatility measure; however, it appears that both $\mathcal{V}\left(F_{X\mid r,\tfrac{m}{m+1}=\tfrac{1}{2}}^{\left(\textrm{NB}\right)}\right)=1.0151$ and $\mathcal{V}\left(F_{X\mid\alpha=1,\vartheta}^{\left(\textrm{War}\right)}\right)=0.9423$ are strong indicators of idiosyncrasy.
+6.3 Pareto 2$\left(\alpha,\vartheta\right)$ vs. Inverse Gamma$\left(r,\lambda\right)$
+In Subsection 6.1, we noted that the 2-parameter Pareto $\textrm{Pareto 2}\left(\alpha,\vartheta\right)$ family possesses a rather low versatility measure compared to the other 2-parameter continuous distributions in Table 3.
+This observation affords the opportunity to consider whether, despite its common use, Pareto $\textrm{Pareto 2}\left(\alpha,\vartheta\right)$ is unduly idiosyncratic.
+Given the distribution’s obvious simplicity, the basic question is whether it is insufficiently adaptable for modeling severities effectively, and therefore should be replaced by other distributions with similar characteristics (in particular, heavy tails) but greater adaptability.
+One potential alternative is the Inverse $\textrm{Inverse Gamma}\left(r,\lambda\right)$ distribution, associated with the random variable $Y$ such that $\dfrac{1}{Y}\sim\textrm{Gamma}\left(r,\lambda\right)$.
+The Inverse $\textrm{Inverse Gamma}\left(r,\lambda\right)$ PDF is given by
+$f_{Y\mid r,\lambda}^{\left(\textrm{I}\Gamma\right)}\left(y\right)=\dfrac{\lambda^{r}e^{-\lambda/y}}{y^{r+1}\Gamma\left(r\right)},$
+where the tail parameter, $r$, plays a role comparable to the Pareto $\textrm{Pareto 2}\left(\alpha,\vartheta\right)$’s $\alpha$ in that for all real $\kappa\leq r$, $\textrm{E}_{Y\mid r,\lambda}^{\left(\textrm{I}\Gamma\right)}\left[Y^{\kappa}\right]=\infty$.
+As illustrated in Figure 2, the PDFs of the two distributions can differ considerably near the lower end of the sample space, with Pareto $\textrm{Pareto 2}\left(\alpha,\vartheta\right)$ possessing a mode at 0 (where $f_{Y\mid\alpha,\vartheta}^{\left(\textrm{P2}\right)}\left(0\right)=\tfrac{\alpha}{\vartheta}>0$) and Inverse $\textrm{Inverse Gamma}\left(r,\lambda\right)$ increasing from $f_{Y\mid r,\lambda}^{\left(\textrm{I}\Gamma\right)}\left(0\right)=0$ until it reaches an internal mode at $\tfrac{\lambda}{r+1}$.
+This greater flexibility, and therefore adaptability, of Inverse $\textrm{Inverse Gamma}\left(r,\lambda\right)$ is reflected in its versatility measure of $\mathcal{V}\left(F_{Y\mid r,\lambda}^{\left(\textrm{I}\Gamma\right)}\right)=3.1264$, which is equal to that of the $\textrm{Gamma}\left(r,\lambda\right)$ distribution and substantially greater than $\mathcal{V}\left(F_{Y\mid\alpha,\vartheta}^{\left(\textrm{P2}\right)}\right)=2.0874$.
+[Uncaptioned image][Uncaptioned image]
+Figure 2. PDFs of Pareto $\textrm{Pareto 2}\left(\alpha=1,\vartheta\right)$ and Inverse $\textrm{Inverse Gamma}\left(r=1,\lambda\right)$ Distributions (from left to right)
+[variable parameter $=0.5$ for dotted line, $1.0$ for solid line, and $2.0$ for dashed line]
+To investigate the relative effectiveness of the Pareto $\textrm{Pareto 2}\left(\alpha,\vartheta\right)$ and Inverse $\textrm{Inverse Gamma}\left(r,\lambda\right)$ models in practice, we perform a series of four simple simulation experiments (I, II, III, and IV) to evaluate each model’s ability to distinguish itself from the other model within a hypothesis-testing framework:
+  • In Experiment I, we take Pareto $\textrm{Pareto 2}\left(\alpha,\vartheta\right)$ as the null hypothesis, generate $10,000$ IID pairs $\left(\alpha_{i}\sim\textrm{Lognormal}\left(\mu=0,\sigma=1\right),\vartheta_{i}\sim\textrm{Lognormal}\left(\mu=0,\sigma=1\right)\right)$, and for each pair generate $n$ IID observations $Y_{j}\sim\textrm{Pareto 2}\left(\alpha_{i},\vartheta_{i}\right)$ (where $n\in\left\{50,100,500,1000,5000,\ldots,100,000\right\}$).
+We then use the $n$ IID $Y_{j}$ to perform a minimum chi-squared goodness-of-fit test (at the $0.05$ level of significance), [^23] In all minimum chi-squared tests, we employ maximum likelihood to estimate parameters and Sturges’ rule to set the number of test bins.
+and record the proportion of false rejections (out of the 10,000 $\left(\alpha_{i},\vartheta_{i}\right)$ pairs) as an estimate of Pr{Type 1 $\Pr\left\{\textrm{Type 1 Error}\right\}$.
+  • In Experiment II, we take Pareto $\textrm{Pareto 2}\left(\alpha,\vartheta\right)$ as the null hypothesis, generate $10,000$ IID pairs $\left(r_{i}\sim\textrm{Lognormal}\left(\mu=0,\sigma=1\right),\lambda_{i}\sim\textrm{Lognormal}\left(\mu=0,\sigma=1\right)\right)$, and for each pair generate $n$ IID observations $Y_{j}\sim\textrm{Inverse Gamma}\left(r_{i},\lambda_{i}\right)$ (where $n\in\left\{50,100,500,1000,5000,\ldots,100,000\right\}$).
+We then use the $n$ IID $Y_{j}$ to perform a minimum chi-squared goodness-of-fit test (at the $0.05$ level), and record the proportion of false retentions (out of the 10,000 $\left(r_{i},\lambda_{i}\right)$ pairs) as an estimate of Pr{Type 2 $\Pr\left\{\textrm{Type 2 Error}\right\}$.
+  • Experiments III and IV are analogous to Experiments I and II, respectively, with the roles of Pareto $\textrm{Pareto 2}\left(\alpha,\vartheta\right)$ and Inverse $\textrm{Inverse Gamma}\left(r,\lambda\right)$ reversed.
+The experimental results in Table 5 show rather clearly that Inverse $\textrm{Inverse Gamma}\left(r,\lambda\right)$ performs better on the whole than Pareto $\textrm{Pareto 2}\left(\alpha,\vartheta\right)$.
+Whereas both models yield satisfactorily low frequencies of Type 1 error (approaching the $0.05$ level of significance from below as $n$ increases), the Pareto $\textrm{Pareto 2}\left(\alpha,\vartheta\right)$ hypothesis is consistently poor in recognizing and rejecting Inverse $\textrm{Inverse Gamma}\left(r,\lambda\right)$ observations (with Type 2 error rates above $0.80$).
+This shortcoming, which does not improve for larger sample sizes, appears to derive from the Pareto $\textrm{Pareto 2}\left(\alpha,\vartheta\right)$’s insufficient sensitivity to variations in the frequency of observations at the lower end of the sample space.
+In short, the model gives too much credence to observations consistent with the hypothesized tail behavior, thereby failing to discern the presence of otherwise non-Pareto $\textrm{Pareto 2}\left(\alpha,\vartheta\right)$ sample characteristics.
+Although too limited to be conclusive, the present analysis provides strong prima facie support for using the proposed versatility measure to identify and eschew the Pareto $\textrm{Pareto 2}\left(\alpha,\vartheta\right)$ model’s idiosyncrasy.
+Table 5. Hypothesis-Test Results from Simulated Data
+I. $\boldsymbol{H_{0}}$: Pareto 2 II. $\boldsymbol{H_{0}}$: Pareto 2 III. $\boldsymbol{H_{0}}$: Inverse Gamma IV. $\boldsymbol{H_{0}}$: Inverse Gamma
+(Pareto 2 Obs.) (Inverse Gamma Obs.) (Inverse Gamma Obs.) (Pareto 2 Obs.)
+Sample Size (n) Pr{Type 1 Error} Pr{Type 2 Error} Pr{Type 1 Error} Pr{Type 2 Error}
+$50$ $0.0000$ $1.0000$ $0.0034$ $1.0000$
+$100$ $0.0056$ $0.8671$ $0.0131$ $0.9999$
+$500$ $0.0181$ $0.8113$ $0.0195$ $0.6822$
+$1000$ $0.0239$ $0.8084$ $0.0252$ $0.5129$
+$5000$ $0.0296$ $0.8153$ $0.0362$ $0.3039$
+$10,000$ $0.0320$ $0.8252$ $0.0379$ $0.2478$
+$50,000$ $0.0402$ $0.8541$ $0.0435$ $0.1599$
+$100,000$ $0.0482$ $0.8645$ $0.0494$ $0.1367$
 7 Conclusion
@@
 These examples showed that:
-  • The versatility measure of a 2-parameter distribution often falls between the measures of two 1-parameter special cases (with each of the parameters fixed in turn), suggesting that the characteristics contributed by the two parameters individually are “averaged” when taken together.
+  • The versatility measure of a 2-parameter distribution often, but not always, falls between the measures of two 1-parameter special cases (with each of the parameters fixed in turn), suggesting that the characteristics contributed by the two parameters individually tend to be “averaged” when taken together.
   • Parameters with a direct effect on tail behavior – for example, $\tau$ for the $\textrm{Weibull}\left(\lambda,\tau\right)$ and Discrete $\textrm{Weibull}\left(\tfrac{m}{m+1},\tau\right)$ distributions and $\alpha$ for the Pareto $\textrm{Pareto 2}\left(\alpha,\vartheta\right)$ and $\textrm{Waring}\left(\alpha,\vartheta\right)$ distributions – appear to have relatively larger impacts on versatility than other parameters.
   • The versatility measures associated with certain discrete families tend to be lower than those of their continuous analogues because versatility is sensitive to the level of functional adaptability afforded in the neighborhood of 00.
-  • The versatility measures of certain distributions (i.e., $\textrm{Lognormal}\left(\ln\left(\nu\right),\sigma=1\right)$, Negative $\textrm{Binomial}\left(r,\tfrac{m}{m+1}=\tfrac{1}{2}\right)$, and $\textrm{Waring}\left(\alpha=1,\vartheta\right)$) are considerably lower than other measures, raising serious concerns about idiosyncratic behavior.
-Overall, we would argue that the proposed versatility measure may be employed to assess the potential usefulness of probability distributions a priori by accounting for and mitigating problems of idiosyncratic overfitting.
-Although our analysis does not propose minimum cutoff values for model-selection purposes, the versatility measure can be used as an effective guide for identifying a range of acceptable values based on the measures arising from distribution families found to be useful in practice.
-For example, the relatively low measures associated with the Pareto $\textrm{Pareto 2}\left(\alpha,\vartheta\right)$ family should not necessarily disqualify this commonly used model, but rather suggest a lower region for reasonable versatility levels.
-On the other hand, the substantially lower measures computed for the $\textrm{Lognormal}\left(\ln\left(\nu\right),\sigma=1\right)$, Negative $\textrm{Negative Binomial}\left(r,\tfrac{m}{m+1}=\tfrac{1}{2}\right)$, and $\textrm{Waring}\left(\alpha=1,\vartheta\right)$ distributions likely suggest these models should be avoided.
-Much further work remains to be done in this area, particularly involving the analysis of marginal contributions to versatility of additional parameters for $k\geq 2$.
-In this regard, it seems likely that the most useful potential applications of the proposed versatility measure lie in the investigation and comparison of probability distributions within multiple-parameter families; for example, the Generalized Beta, Burr, Pareto, and Pearson families increasingly used to model severities in risk analysis.
+  • The versatility measures of certain distributions (i.e., Pareto $\textrm{Pareto 2}\left(\alpha,\vartheta\right)$, $\textrm{Lognormal}\left(\ln\left(\nu\right),\sigma=1\right)$, Negative $\textrm{Negative Binomial}\left(r,\tfrac{m}{m+1}=\tfrac{1}{2}\right)$, and $\textrm{Waring}\left(\alpha=1,\vartheta\right)$) are noticeably lower than those of other distributions, raising serious concerns about idiosyncratic behavior.
+Finally, by comparing the Pareto $\textrm{Pareto 2}\left(\alpha,\vartheta\right)$ model with the more versatile Inverse $\textrm{Inverse Gamma}\left(r,\lambda\right)$ alternative, we showed how the proposed versatility measure may be employed to assess the usefulness of probability distributions a priori, effectively recognizing and mitigating problems of idiosyncratic overfitting.
+Although our analysis does not propose minimum cutoff values for model-selection purposes, we believe that, by observing large numbers of versatility measures in applied work, it will be possible to establish reasonable empirical guidelines for acceptable ranges of $\mathcal{V}\left(F_{0}\right)$.
+Much further research remains to be done on this topic, particularly involving the analysis of marginal contributions to versatility of additional parameters for $k\geq 2$.
+In this regard, it seems likely that the most useful potential applications of the proposed versatility measure lie in the investigation and comparison of probability distributions within: (i) multiple-parameter continuous families, such as the Generalized Beta, Burr, Pareto, and Pearson distributions used to model severities; and (ii) discrete families modified to fit individual parameters to specific values within the sample space (especially $x=0$ or $1$), such as the Zero-Inflated Poisson and Negative Binomial models for frequencies.